### v1

# Token Reduction Should Go Beyond Efficiency in Generative Models – From Vision, Language to Multimodality

Abstract
In Transformer architectures, tokens—discrete units derived from raw data—are formed by segmenting inputs into fixed-length chunks. Each token is then mapped to an embedding, enabling parallel attention computations while preserving the input’s essential information. Due to the quadratic computational complexity of transformer self-attention mechanisms, token reduction has primarily been used as an efficiency strategy. This is especially true in single vision and language domains, where it helps balance computational costs, memory usage, and inference latency. Despite these advances, this paper argues that token reduction should transcend its traditional efficiency-oriented role in the era of large generative models. Instead, we position it as a fundamental principle in generative modeling, critically influencing both model architecture and broader applications. Specifically, we contend that across vision, language, and multimodal systems, token reduction can: (i) facilitate deeper multimodal integration and alignment, (ii) mitigate "overthinking" and hallucinations, (iii) maintain coherence over long inputs, and (iv) enhance training stability, etc. We reframe token reduction as more than an efficiency measure. By doing so, we outline promising future directions, including algorithm design, reinforcement learning-guided token reduction, token optimization for in-context learning, and broader ML and scientific domains. We highlight its potential to drive new model architectures and learning strategies that improve robustness, increase interpretability, and better align with the objectives of generative modeling.  [^1] We collected a list of token reduction papers at Github.
1 Introduction
Transformer-based generative models brown2020language ; devlin2019bert ; NEURIPS2021_854d9fca ; vaswani2017attention have emerged as dominant deep learning architectures across vision, language, and multimodal tasks, due to their ability to process long sequences of tokens, which are the fundamental representational units derived from raw data such as subwords in language or image patches in vision. As these models are applied to increasingly complex real-world tasks, the input sequence lengths of both the models and their training datasets continue to grow. However, the quadratic computational complexity of the attention mechanism results in high memory usage and slow inference, which hinders the practical deployment of generative models at scale. Token reduction addresses this challenge by reducing the number of tokens processed during inference. By pruning or merging tokens, token reduction decreases computational cost and accelerates runtime, offering a practical solution for improving the efficiency of generative models.
Token reduction has been widely adopted in computer vision, language processing, and multimodal tasks. In vision transformers, it has primarily been used to reduce computational cost by removing visually redundant tokens bolya2022token ; kong2022spvit ; liang2022not ; rao2021dynamicvit . While effective in decreasing token-level computations and accelerating inference, this approach can discard subtle but important visual features, especially in dense prediction tasks such as segmentation and object detection, leading to significant performance degradation. In language models, token reduction has commonly been implemented through early-exit mechanisms and token-skipping strategies lin2025boosting ; wu2024accelerating , which reduce the number of intermediate tokens processed and thus lower computational overhead. Similarly, multimodal large language models (MLLMs) apply visual token pruning primarily during the prefill stage chen2024image , where adaptive attention patterns are learned in the early layers to prune tokens in later stages. Despite progress in existing work, token reduction is still predominantly viewed as a means of improving computational efficiency, primarily by reducing the number of tokens to minimize associated computations and accelerate inference. Such an efficiency-only mindset has critical limitations. Naive pruning methods may discard informative tokens, thereby degrading model understanding and performance liang2022not ; zhan2024exploring ; zhan-etal-2024-rethinking-token . Furthermore, token reduction is commonly treated as a post hoc optimization, rather than being integrated into the core design and training of the model chen2024image .
In this paper, we argue that viewing token reduction purely from an efficiency perspective is fundamentally limited. Instead, we position token reduction as a core design principle in generative modeling, deeply integrated with both training and inference to prioritize tokens that maximize downstream task performance and semantic integrity.
Modern generative tasks present numerous challenges that highlight the need for thoughtful token selection: (i) Ultra-long contexts in language modeling require selective retention of relevant segments to preserve coherence. (ii) LLMs frequently exhibit overthinking, repeatedly attending to low-value tokens and producing redundant or contradictory outputs. (iii) Multimodal generation tasks often face issues of visual redundancy, where background tokens overshadow salient visual features critical for accurate understanding. (iv) Noisy or irrelevant tokens introduced during training slow down convergence and harm model stability. By learning to intelligently select, merge, or compress tokens based on their contribution to generation objectives-rather than solely on raw redundancy-models can simultaneously reduce computational load, improve robustness, and enhance interpretability and alignment. This paper makes the following three key contributions:
  • We provide a comprehensive analysis of the limitations inherent in current efficiency-centric token pruning methods across vision, language, and multimodal generative models, highlighting critical shortcomings in these widely adopted strategies.
  • We identify core challenges faced by modern generative models including insufficient visual representation, semantic misalignment, overthinking in reasoning, and training instability. We then demonstrate how principled token reduction strategies can effectively mitigate these issues.
  • We propose a roadmap for future research on token reduction, including directions for method design, reinforcement learning-guided token selection, adaptive in-context compression, and hardware-algorithm co-design, etc. These directions aim to support the development of next-generation generative architectures that are both robust and efficient.
This position paper is organized as follows: Sec. 2 reviews prior token reduction methods across various modalities. Sec. 3 introduces the problem formulation, Sec. 4 formalizes the identified challenges and demonstrates how informed token reduction strategies can address them. Sec. 5 proposes promising research directions for advancing token reduction as well as broader implications.
2 Related Work
2.1 Token Reduction in Vision Models
Image Classification. Classification serves as a fundamental task for vision models and token reduction techniques have been widely applied in it due to its simplicity and versatility. It has been widely explored from various aspects bolya2022token ; liang2022not ; rao2021dynamicvit ; wu2023ppt ; zeng2024m2m . Specifically, DynamicViT rao2021dynamicvit devises a lightweight module to predict the importance score of each token, thereby pruning unimportant tokens. E-ViT liang2022not identifies attentive tokens from feed-forward network, enabling token pruning without additional parameters. ToMe bolya2022token merges tokens with similarity based on bipartite matching to maintain information utility. PPT wu2023ppt analyzes the statistic data between layers and adaptively employs token pruning and merging within layers to achieve higher acceleration performance.
Year2020 - 2021Power-bert goyal2020power , DynamicViT rao2021dynamicvit ,TokenLearner ryoo2021tokenlearner , SpAtten wang2021spatten 2022LTP kim2022learned , SPViT kong2022spvit , EViT liang2022not ,PatchSlim tang2022patch , STTS wang2022efficient , PPT wu2023ppt 2023DCP anagnostidis2023dynamic , ToMe bolya2022token , ToMeSD bolya2023token , PuMer cao2023pumer ,HeatViT dong2023heatvit , Llmlingua jiang2023llmlingua ,DTP nawrot2022efficient , TPS wei2023joint 2024Honeybee cha2024honeybee , FastV chen2024image , TokenPacker li2024tokenpacker ,BRAVE kar2024brave , LLaMA-VID li2024llama , DyDiT zhao2024dynamic - 2025.05DivPrune alvar2025divprune , SeTok wu2024towards , TopV yang2025topv ,Voco-llama ye2024voco , S2-MAD zeng2025s , ToCa zou2025accelerating Vision & Language TransformerLLM & Multimodal LLMEfficiency
$\bm{\checkmark}$
$\bm{\Rightarrow}$  Future
Figure 1: Timeline of notable method developments for token reduction methods with modality shifts (Vision & Language $\to$ Multimodal LLMs). All these strategies aim to speed up inference with neglected performance drops. Conversely, we ask: What is the next token reduction paradigm in generative model design that goes beyond test-time accelerations?
Video Compression. Unlike token reduction in image classification, video compression focuses more on the temporal redundancy within videos, and algorithms are developed to reduce the number of tokens with less computational overhead. Various token reduction methods have been investigated for different tasks including video understanding ryoo2021tokenlearner ; wang2022efficient , video editing li2024vidtome , video-text retrieval liu2022ts2 ; shen2024tempme , video action detection chen2023efficient and so on. Specifically, STTS wang2022efficient introduces a lightweight framework that dynamically selects the most informative spatial-temporal tokens in video transformers. Tokenlearner ryoo2021tokenlearner proposes an adaptive tokenization module that learns a handful of informative spatial-temporal tokens, significantly reducing computational costs. EVAD chen2023efficient selectively drops irrelevant spatial-temporal tokens in non-keyframes while preserving keyframe and motion-relevant tokens, and then refines actor features using a context-aware decoder to maintain accuracy with reduced computations.
Generative Tasks. Token reduction in generative tasks ju2024turbo aims to accelerate generative models through the efficient utilization of tokens. It can be applied to both diffusion models bolya2023token ; wang2024attention and diffusion transformers zou2025accelerating . Specifically, ToMeSD bolya2023token exploits natural redundancy in generated images by merging redundant tokens, successfully extending token merging to stable diffusion with simple unmerging. DyDiT zhao2024dynamic reduces redundancy with a Timestep-wise Dynamic Width approach to adopt model width conditioned on the generation timesteps, and a Spatial-wise Dynamic Token strategy to avoid redundant computations at unnecessary spatial locations.
2.2 Token Reduction in Language Models
Token reduction strategies in language modeling have evolved from early optimizations for BERT huang2022pyramid ; kim2020length ; kim2022learned ; kim2023leap ; ye2021tr to techniques specifically designed for LLMs. PoWER-BERT goyal2020power introduces progressive word-vector elimination by removing redundant token representations based on self-attention dynamics, improving inference efficiency. Learned token pruning kim2022learned extends this approach by learning attention-based thresholds to adaptively prune uninformative tokens, thereby reducing computational costs while preserving model performance. In LLMs, token reduction must account for the constraints of autoregressive decoding across diverse downstream tasks. Dynamic pooling methods anagnostidis2023dynamic ; tao2025saliency adjust token representations on the fly during inference to reduce redundancy. Prompt compression techniques fu2024lazyllm ; ge2023context ; jiang2023llmlingua aim to reduce computational overhead by compressing the input prompt before generation. Selective decoding approaches fu2024lazyllm ; wingate2022prompt reduce per-step inference costs by computing key-value pairs only for tokens critical to predicting the next token. In multi-agent systems, S2-MAD zeng2025s proposes a sparsification mechanism that limits unnecessary token exchanges between agents, reducing communication costs and improving the efficiency of collaborative reasoning.
2.3 Token Reduction in Multimodal LLMs
Recent work has explored visual token pruning by addressing attention inefficiencies of deep transformer layers in MLLMs alvar2025divprune ; arif2025hired ; lin2025boosting ; shang2024llava . Specifically, FastV chen2024image shows that deeper vision-language layers expend significant computations on redundant image tokens. To address this, a lightweight module is adopted to adaptively prune these tokens, reducing inference overheads in subsequent stages. A complementary approach modifies vision feature extractors or projectors to output a smaller set of highly informative image tokens, effectively distilling the input into a compressed representation cai2024matryoshka ; cha2024honeybee ; kar2024brave ; li2024tokenpacker ; li2024mini ; ye2024voco . However, efficiency gains from these prefill stages often fade during the decoding phase, where per-token computations dominate. To overcome this, recent methods jointly optimize token reduction during both prefill and decoding stages, ensuring sustained speedups throughout inference huang2024dynamic ; song2024less . In Fig. 1, we present a timeline of notable developments in token reduction methods, illustrating the shift from early applications in ViT and BERT-based models to more recent advances in LLMs and MLLMs.
3 Problem Formulation
In modern generative models Qwen2.5VL ; 2024llama3 ; liu2023llava ; openai2024gpt4 ; Peebles2022DiT , a token denotes one fundamental unit of input or representation, typically encoded as a vector. For example, a token might correspond to a subword in language, a patch in an image, or an embedding of a time step in audio. We denote a sequence of $N$ input tokens as $X=[x_{1},\dots,x_{N}]\in\mathbb{R}^{d}$. Token reduction refers to any operation that compresses the token sequence to $M$ tokens (with $M<N$) by removing or consolidating tokens while aiming to preserve the original information.
Broadly, token reduction methods fall into the following categories: 1) Token pruning methods bolya2022token that remove entire unimportant tokens, simply dropping them from the sequence; 2) Token merging methods bolya2022token ; bolya2023token which fuse information from multiple tokens into fewer tokens, effectively compressing the sequence by merging similar or related tokens; 3) Hybrid strategies cao2023pumer ; kim2024token ; wu2023ppt that combine pruning and merging within a unified framework; 4) Token distillation approaches cai2024matryoshka ; mu2023learning which integrate rich information across longer input sequences or multiple modalities into fewer condensed tokens, enabling efficient cross-modal interactions and long-context reasoning in LLMs and MLLMs.
A core challenge in token reduction is the determination of tokens to be pruned or merged. There are various importance criteria and scoring mechanisms to rank token significance, including attention-based heuristics liang2022not , gradient or loss-based criteria huang2024dynamic , clustering haurum2024agglomerative , and learned predictors rao2021dynamicvit .
From a purely efficiency-oriented perspective, token reduction delivers substantial computational efficiency gains by reducing the quadratic computation cost from $\mathcal{O}(N^{2})$ to $\mathcal{O}(M^{2})$ in attention mechanisms. By eliminating redundant tokens and processing fewer computations during inference, it effectively accelerates the inference speed and improves the model throughput, which is crucial for latency-sensitive tasks or real-time applications. Furthermore, it also shrinks the memory footprint for activations and gradients (e.g., key/value caches), alleviating memory usage for both inference and training, which is particularly beneficial for wide deployments on resource-limited platforms.
However, as stated in this position paper, token reduction can benefit models in multiple ways beyond efficiency, which will be introduced in detail in the following sections.
4 Challenges and Positions
In this section, we establish token reduction as a foundational mechanism for addressing critical challenges in modern generative systems. We claim five core challenges across modalities: visual representation sparsity, semantic misalignment, reasoning redundancy, training instability, and long-context overload. We demonstrate how principled token reduction strategies intrinsically address these issues through dynamic token-semantic co-optimization. We position token reduction not only as an efficiency tool, but as an essential paradigm for enhancing semantic coherence and enabling sustainable scaling of generative systems.
4.1 Obtain Informative Visual Representation
MLLMs often suffer from noisy visual inputs that impede fine-grained understanding. We outline key challenges in MLLM visual reasoning: ① Text-Visual Attention Shift: Due to the rotary positional embeddings in LLM decoders, later text tokens disproportionately attend to spatially lower image regions hong2024token , shifting attention away from semantically important areas (e.g., objects at the top of an image); ② Visual Redundancy: Empirical studies lin2025boosting ; wu2024accelerating show that beyond the first few layers, many image tokens contribute little new information, ③ Task-Guided Focus in VQA: In multimodal question answering, the question itself pinpoints relevant image regions (e.g., "kitten color" directs focus to the kitten patch), implying that many image tokens are unnecessary for correct answers song2024less .
Therefore, we position token reduction as a representation-learning optimization: selecting the subset of tokens that preserves informative visual representation. For example, VisPruner zhang2025textvisualattentionexploitingvisual identifies high-value tokens using visual-encoder attention and removes duplicates via clustering to ensure diversity. VTW lin2025boosting observes that visual information migrates into text tokens within early layers; it therefore withdraws all visual tokens after a chosen layer based on KL-divergence criteria. TRIM song2024less leverages the CLIP metric and IQR scoring function to adaptively select image tokens that are crucial for answering questions, while an aggregated token is used to retain additional image information.
4.2 Better Multimodal Token Alignment
Despite their impressive capabilities, MLLMs continue to face challenges in semantic alignment. Standard vision tokenizers typically split images into fixed-size patches, which can fragment coherent visual entities (e.g., objects or regions) across multiple tokens. This fragmentation weakens the alignment between visual and linguistic representations. Token reduction offers a promising solution by selecting visual tokens based on semantic importance, thereby producing a compact set of tokens that better align with language representations. Specifically, SeTok wu2024towards dynamically clusters visual features into semantically meaningful tokens using a density-peak algorithm, which determines both the number and structure of token groupings per image. This approach preserves both high- and low-frequency semantics, substantially improving concept-level alignment and downstream task performance. M3 cai2024matryoshka introduces a hierarchical token structure that captures coarse-to-fine semantic granularity, allowing different levels of abstraction to be selectively retained depending on task needs.
4.3 Reduce Overthinking in Reasoning
LLM reasoning. In the context of language models, overthinking refers to generating excessively long or convoluted chains of reasoning that go beyond what is necessary to reach a correct answer. An LLM may produce verbose, repetitive, or even self-contradictory explanations when it fails to converge on a solution-often due to uncertainty sui2025stop ; wang2025harnessing . Such extended reasoning trajectories are inefficient and recent studies show that state-of-the-art reasoners can consume over 15,000 tokens to solve math problems that could be addressed with a concise chain-of-thought (CoT) of just a few hundred tokens hou2025thinkprune . Mitigating overthinking is thus crucial for improving both accuracy and efficiency in generative modeling. By reducing unnecessary tokens during reasoning, LLMs can focus on salient steps, aligning generation with a more concise and logical trajectory.
CoT-Influx huang2023fewer introduces a CoT pruning strategy in which concise reasoning examples are included in the prompt. By pruning unimportant tokens from these examples, more reasoning demonstrations can fit into the context window, surprisingly leading to improved math reasoning accuracy. TokenSkip xia2025tokenskip enables LLMs to skip less important tokens within CoT sequences and learn shortcuts between critical reasoning steps. This allows for controllable CoT compression with adjustable compression ratios, enabling models to automatically trim redundant tokens during reasoning.
MLLM reasoning. MLLMs, which reason over text and other modalities, face similar overthinking issues. In vision-language tasks, overthinking often manifests as excessive processing of visual tokens or overly detailed image descriptions, resulting in inefficiency and potential confusion chen2025zipr1reinforcingtokensparsity . Token reduction techniques in MLLMs aim to promote more focused and sparse reasoning over multimodal inputs. For example, FAST xiao2025fast rewards shorter-than-average token sequences for correct answers, while allowing longer reasoning for more complex tasks. It also adjusts policy optimization constraints to tighten output exploration for simple tasks (thus reducing unnecessary tokens) and loosen it for harder ones to allow deeper reasoning.
Together, these strategies reduce overthinking in straightforward cases, boosting efficiency while preserving effective reasoning depth for complex scenarios.
4.4 Improve Training Stability
While token reduction has traditionally been employed as a post-training optimization to enhance inference efficiency, recent research indicates its potential to significantly improve training stability when integrated into the pre-training phase gao2023masked ; li2024pruning ; lin2024not , suggesting that selective token utilization during training can lead to more robust and effective model learning.
One notable approach is Rho-1 lin2024not , which involves scoring tokens based on their alignment with a desired distribution using a reference model and then focusing the training loss on tokens with higher scores. Therefore, it effectively filters out noisy or less informative tokens, leading to faster convergence and improved performance. UPFT ji2025first emphasizes the importance of initial reasoning steps in training. By reducing the number of training tokens, UPFT encourages the model to focus on the initial prefix substrings of reasoning trajectories, which are often more stable and contain crucial information. This focus helps the model avoid being influenced by subsequent complex or potentially erroneous information, thereby improving training stability. However, how to design token reduction algorithms that are tightly integrated with training procedures such as in modern methods like GRPO shao2024deepseekmath , remains underexplored. Future research should investigate specialized approaches that incorporate token reduction directly into training objectives, enabling models to learn to prioritize or discard tokens in a task-aware and gradient-aligned manner.
4.5 Enhance Long Context & Video Understanding
Long-context LLMs. Long-context language modeling presents unique challenges: ① Long texts often contain raw tokens that exhibit repetitive descriptions and irrelevant details that strain the attention mechanism; ② LLM-based agent systems use input data as sequential prompts for reasoning or for switching between multiple tasks, which can lead to overload when the prompt grows too large; ③ It is very difficult to scale up to even longer content for learning more information. Token reduction techniques directly address these issues by distilling extensive input sequences into compact summary vectors or representative tokens. By doing so, models preserve core information such as key events, central themes, or task-specific facts, while significantly decreasing cognitive load. For example, AutoCompressors chevalier2023adapting trains pre-trained LLMs to compress long contexts into compact summary tokens, reducing token length by orders of magnitude to extend context windows and speed up inference. TokenSwift wu2025hours reduces the effective number of tokens that the model dynamically processes during generation by using multi-token parallel generation and n-gram retrieval for token reutilization, therefore enabling efficient ultra-long sequence generation (up to 100K tokens).
Video-based MLLMs. The necessity of token reduction primarily lies in enhancing the model’s effective understanding of video content through: ① Instruction-guided information filtering: token reduction prioritizes selecting visual information relevant to user instructions over raw data volume. ② preserving spatiotemporal structure: token reduction strategically compresses massive spatiotemporal information to retain spatiotemporal dependencies, ensuring the model can capture dynamic semantics, as well as prevent redundant tokens interfere with long temporal reasoning. ③ Preserving semantic integrity: it facilitates feasible processing of extremely long sequences in learning while preserving semantic integrity. ④ Multi-modal alignment: token reduction distills visual information into a compact, semantically aligned form, thereby efficiently bridging the gap between language and vision liu2025hybrid . By doing so, it effectively addresses the challenges posed by the low abstractness and lack of guidance inherent in raw visual inputs, which are the root causes of semantic misalignment and optimization ambiguity in multi-modal models. Recent works illustrate these principles: HICom liu2025hybrid conducts conditional token compression at local and global levels using user instructions as guidance to retain instruction-relevant visual information while reducing computational burden. Video-XL-Pro liu2025video employs reconstructive token compression with a dynamic token synthesizer and semantic-guided masking to generate compact yet comprehensive video tokens for improved MLLM performance and efficiency.
5 Future Directions
In this section, we propose eight promising directions for token reduction beyond the efficiency benefits, organized into three categories: (i) Algorithmic Innovations (Sec. 5.1$\sim$5.4), (ii) Application Innovations (Sec. 5.5$\sim$ 5.7), and (iii) Hardware-Algorithm Co-Design (Sec. 5.8).
5.1 Design of New Algorithms
Future research on algorithm design should explore holistic and adaptive token reduction strategies. Building on recent advances, we outline six promising directions:
Better Token Importance Metrics. It is critical to re-evaluate how token importance is defined and measured. More robust and unbiased scoring mechanisms can be developed, such as predictors akhauri2025tokenbutler or meta-learning frameworks that go beyond attention-based proxies. These models should capture downstream utility with minimal supervision, enabling adaptive pruning across tasks and domains.
Constructive Token Compression. Token reduction can shift from purely eliminative pruning to strategies that merge spatially or semantically similar tokens into compact summary vectors li2024tokenpacker .
Mitigating Position Bias. In MLLMs, attention-based pruning methods (e.g., FastV) often rely on attention scores from a fixed query token, leading to retained tokens concentrating in specific image regions (e.g., lower corner) wen2025token with potential position bias. Future methods should preserve spatial diversity by enforcing structural uniformity in retained tokens to improve robustness on visual tasks.
Cross-Modal Guided Pruning. Pruning decisions in MLLMs should be guided by inter-modality dependencies, rather than made independently for each modality. For example, text-guided pruning of visual tokens can improve alignment between modalities cao2024madtp . The design should account for joint representations and semantic correspondence across all relevant inputs.
End-to-End Sparsification. Token reduction should consider both the prefill stage and decoding phase for LLMs. This includes dynamically managing the sparsity of KV caches and selectively updating generated tokens, sustaining efficiency gains throughout the entire inference process  huang2024dynamic .
Hardware-Algorithm Co-Design. Token pruning can explore custom hardware and compiler optimizations that take advantage of dynamic token sparsity patterns (e.g., irregular memory access and conditional computation) to maximize throughput and energy efficiency as detailed in Sec. 5.8.
5.2 Token Reduction for Reinforcement Learning
Reinforcement learning (RL)-driven token reduction technology demonstrates significant potential in multimodal large models, with its core lying in balancing computational efficiency and reasoning accuracy through dynamic reward mechanisms and sparsity constraints. Under the “Fast-Slow Thinking” framework xiao2025fast , RL can guide hierarchical screening of high-value tokens: The fast path employs sparsity strategies (e.g., rule-based rewards or information density scoring) to filter redundant visual or semantic features, while the slow path focuses on refined reasoning. This hierarchical mechanism not only reduces computational overhead by up to 30-50%, but also continuously optimizes token selection strategies via RL’s online feedback. For instance, integrating self-play mechanisms with the GRPO algorithm (Group Relative Policy Optimization) enables the generation of diverse candidate paths and selects optimal outputs.
Furthermore, the sparsity reinforcement paradigm proposed in ZipR1 chen2025zipr1reinforcingtokensparsity highlights RL’s unique advantage in cross-modal generalization. By designing sparsity rewards (e.g., token quantity limits or step compression incentives), models can autonomously generate concise intermediate reasoning chains and even distill such capabilities into lightweight models. For example, in visual question answering tasks, models can dynamically extract key image region features while enhancing answer reliability through multi-candidate contrast mechanisms (e.g., parallel samplings), achieving the goal of “low computational cost, high task accuracy.” Looking ahead, this technology could advance toward adaptive sparsity and cross-modal alignment, providing efficient inference foundations for edge computing and real-time interaction scenarios. Simultaneously, it will drive MLLMs toward co-evolution of lightweight design and strong robustness.
5.3 From Prompt Tuning to In-context-learning
Current token reduction efforts for prompts have primarily aimed at compressing prompts for efficiency, often with impressive results jiang2023llmlingua ; mu2023learning . Looking forward, token reduction should evolve into enhancing reasoning and maximizing utility per token in context. Instead of focusing solely on making prompts shorter, future research should explore how each remaining token can carry more information or trigger more complex inference during in-context learning. One direction is to alter the generation paradigm itself, for example, training language models to predict multiple tokens per step gloeckle2024better . Another idea is to enable deeper internal reasoning without increasing prompt length. CoD adams2023sparse introduces an iterative approach using token reduction within fixed-length constraints: Each step incorporates new salient entities into the summary by compressing/abstracting existing content, forcing the model to enhance density without expanding length.
In summary, the next phase of token reduction research should shift focus from simple prompt compression to reasoning-centric compression. Rather than just trimming prompts, we should ask: How can we make each token in the prompt or context do more work for us? This involves training models with objectives that reward higher-level inference per token, developing architectures that recycle tokens for multi-step thinking, or dynamically selecting the most salient tokens to keep at each stage of dialogue.
5.4 Complementary to Other Methods
Token reduction can complement other efficiency techniques, such as quantization. By selectively reducing the number of tokens processed during inference, models can improve both performance and efficiency, particularly when paired with quantization strategies li2025mergevq . Traditional key-value cache quantization methods often suffer from accuracy loss due to their inability to handle outlier tokens that carry distinct or rare features. To mitigate this issue, Outlier Token Tracking su2025accuratekvcachequantization identifies outlier tokens during decoding and excludes them from quantization, preserving full-precision necessary representations and improving key-value cache quantization accuracy. Similarly, Agile-Quant shen2024agile incorporates token pruning as a preprocessing step to reduce the impact of activation outliers. It prunes tokens based on their attention to the start-of-sequence token, discarding those with low attentiveness, which often appear in adjacent channels and contribute to quantization noise. This targeted pruning reduces interaction distances between salient tokens and helps maintain model accuracy under low-bit quantization settings.
5.5 Towards Dense Prediction Tasks for Vision
Existing works primarily concentrate on compressing the backbone of models to ensure their generalization ability, and few works explore recovering all tokens for dense prediction tasks bolya2023token ; liang2022expediting . It is necessary to develop custom token reduction methods for various downstream dense prediction applications like autonomous driving and robotic control with specific settings and requirements. Lacking these specialized designs would lead to a mismatch and performance drop when deployed in real-world settings. For example, autonomous driving zhang2024sparsead would require displacement and velocity based on occupancy prediction, and robotic control wang2024sparse would demand rotation angle according to the grid map. Therefore, how to develop fast and specialized token reduction strategies tailored for downstream dense prediction tasks is crucial for deployment in practical scenarios.
5.6 Towards Long Video Applications
Exploiting long videos holds great potential, as processing hours of footage is significantly more labor-intensive and time-consuming than working with short clips. Due to the inherent complexity and resource demands, most current research on long video learning focuses on discriminative tasks such as video understanding lee2024video ; ren2023testa . In contrast, broader applications including long video editing zhang2025adaflow , long video-text retrieval, and narrative-level generation remain largely underexplored. Progress in these areas could have a significant impact on scene editing in video clips, character rendering in movies, and retrieving useful information from numerous videos. Moreover, token reduction offers a path toward interpretability and efficiency in long video processing jiang2025token . This mimics the human visual system, which does not attend to every frame in detail but instead focuses on salient spatiotemporal changes, such as actions or object movement, while filtering out static, redundant content like backgrounds and stationary objects. Future models should similarly prioritize informative frames and temporal segments, allowing them to reason over extended video sequences with greater efficiency and interpretability.
5.7 Towards AI for Broader ML and Scientific Domains
Token reduction methods can also offer powerful opportunities to reshape broader machine learning and scientific applications. In particular, domains such as medicine, biology, chemistry, and temporal data analysis frequently encounter complex data structures, heterogeneous data sources, and intricate domain-specific relationships. Informed tokenization approaches promise to address these challenges by transforming complex and rich scientific data into concise, informative, and flexible representations, significantly enhancing the utility of transformer-based foundation models across these domains.
Building Biomedical Tokenizers. Recent works exemplify the transformative potential of advanced tokenization methods in the biomedical domain, including protein yuan2025protein ; gao2025foldtoken ; suyunu2025evobpe , genomic eapen2025genomic , and chemical structure yan2024invariant tokenizers. Collectively, these methods illustrate how informed reduction and condensation of input tokens can lead to more effective and interpretable scientific models. For example, traditional tokenizers in EHR foundation models typically treat medical codes as isolated textual units, neglecting their inherent structured and relational context, such as hierarchical relationships, disease co-occurrences, and drug-treatment associations found within biomedical ontologies. To solve this issue, MedTok su2025multimodal integrates textual descriptions and graph-based relational data into a unified tokenization framework. It first uses a language model encoder to extract embeddings from medical code descriptions and employs a graph encoder to capture relational structures from biomedical ontologies. These embeddings are combined into a compact token space through vector quantization, preserving both modality-specific and cross-modality information.
To enhance informativeness and reduce redundancy, MedTok employs a token packing mechanism. It optimizes shared tokens and modality-specific tokens, ensuring that the final tokens encode both shared semantic meaning and modality-specific structure. This process drastically reduces effective vocabulary size, addressing the scalability challenge of 600,000+ medical codes by collapsing redundant representations while preserving critical clinical context. Inspired by adaptive tokenization methods for vision duggal2024adaptive ; yan2024elastictok ; kang2023tictok , future EHR tokenization would be adaptive, enabling dynamic representation of patients’ medical histories, where the length of the token series of each patient history would be directly correlated to length and complexity. Such adaptive tokenization can significantly improve training and inference efficiency across diverse healthcare systems.
Time-Series Data and Clinical Reasoning. Temporal dynamics form an essential component of clinical reasoning, particularly through longitudinal patient data like lab tests and vital signs. However, current large language models struggle to effectively incorporate time-series inputs due to challenges in temporal tokenization spathis2024first ; anjum2024lipcot ; fang2025tsla ; spathis2024first ; masserano2024enhancing . Future tokenization methods should not only dynamically adjust the number of tokens according to temporal complexity but also selectively focus on time segments most relevant to the clinical context, prompt, or task at hand talukder2024totem . This could enhance training effectiveness and inference accuracy, helping create the next generation of EHR foundation models, which are flexible not only over different tasks or prompts, but also over different data sources, patients, and populations. The complexity and richness of EHR data offer opportunities for AI-driven advancements in patient health outcomes. Future EHR models should support comprehensive reasoning capabilities, encompassing complete patient histories, such as vitals, lab results, diagnoses, and procedures over time. They could facilitate timely disease predictions, accurately forecast chronic disease trajectories, and anticipate patient responses to treatments.
5.8 Algorithm-Hardware Co-Design
While algorithmic advancements in token reduction have achieved impressive computational savings, the next crucial step is to integrate these techniques with hardware-aware design principles. We posit that algorithm-hardware co-design is essential for holistic optimization across the compute stack, considering the interplay between algorithmic choices, hardware architectures (specialized data paths, memory hierarchies, communication fabrics, control logic, etc.), and compiler/runtime support (efficient sparse mapping, dynamic scheduling, irregular-data management, etc.) dong2023heatvit ; parikh2024accelerating .
Currently, co-design efforts targeted at token reduction lag significantly behind pure algorithmic research. This gap is problematic because hardware design needs to balance PPA (power, performance, and area), platform specifics, data movement costs, control overhead, and scalability/reusability 10.1109/ASP-DAC58780.2024.10473968 . Algorithms developed in isolation often generate sparse or irregular compute patterns that general-purpose hardware cannot exploit effectively. Therefore, future research should aim to: 1) Design parameterizable, reconfigurable accelerator modules-such as on-the-fly importance-scoring units and sparse-data pipelines-that natively support token-reduced Transformers. 2) Explore Processing-in-Memory (PIM) architectures to alleviate severe memory bottlenecks caused by dynamic token pruning. By executing scoring operations or partial attention mechanisms within or near memory arrays, PIM can drastically reduce data movement costs and improve end-to-end efficiency.
6 Conclusion
In this position paper, we have argued that token reduction must evolve beyond a mere efficiency optimization to become a core design principle in generative modeling. We have shown how principled token reduction can address key challenges such as enhancing semantic fidelity in vision-language alignment, curbing verbose reasoning trajectories, preserving long-range coherence, and stabilizing learning dynamics. Looking ahead, we outlined a roadmap of promising directions. We anticipate that embracing token reduction as a holistic, task-aware mechanism will yield the next generation of generative architectures that are not only more efficient but also more robust, interpretable, and aligned with real-world demands.

### v2

@@
 We highlight its potential to drive new model architectures and learning strategies that improve robustness, increase interpretability, and better align with the objectives of generative modeling.
-[^1] We collected a list of token reduction papers at Github.
+[^1] We collected a list of token reduction papers at: Awesome-Collection-Token-Reduction.
 1 Introduction
@@
 Classification serves as a fundamental task for vision models and token reduction techniques have been widely applied in it due to its simplicity and versatility.
-It has been widely explored from various aspects bolya2022token ; liang2022not ; rao2021dynamicvit ; wu2023ppt ; zeng2024m2m .
+It has been widely explored from various aspects bolya2022token ; liang2022not ; rao2021dynamicvit ; wu2023ppt ; zeng2024m2m ; kong2023peeling ; zeng2025token .
 Specifically, DynamicViT rao2021dynamicvit devises a lightweight module to predict the importance score of each token, thereby pruning unimportant tokens.
-E-ViT liang2022not identifies attentive tokens from feed-forward network, enabling token pruning without additional parameters.
+SPViT kong2022spvit introduces a soft pruning technique, which integrates the less informative tokens generated by the selector module into a package token that will participate in subsequent calculations rather than being completely discarded.
+EViT liang2022not identifies attentive tokens from the attention map, enabling token pruning without additional parameters.
 ToMe bolya2022token merges tokens with similarity based on bipartite matching to maintain information utility.
 PPT wu2023ppt analyzes the statistic data between layers and adaptively employs token pruning and merging within layers to achieve higher acceleration performance.
-Year2020 - 2021Power-bert goyal2020power , DynamicViT rao2021dynamicvit ,TokenLearner ryoo2021tokenlearner , SpAtten wang2021spatten 2022LTP kim2022learned , SPViT kong2022spvit , EViT liang2022not ,PatchSlim tang2022patch , STTS wang2022efficient , PPT wu2023ppt 2023DCP anagnostidis2023dynamic , ToMe bolya2022token , ToMeSD bolya2023token , PuMer cao2023pumer ,HeatViT dong2023heatvit , Llmlingua jiang2023llmlingua ,DTP nawrot2022efficient , TPS wei2023joint 2024Honeybee cha2024honeybee , FastV chen2024image , TokenPacker li2024tokenpacker ,BRAVE kar2024brave , LLaMA-VID li2024llama , DyDiT zhao2024dynamic - 2025.05DivPrune alvar2025divprune , SeTok wu2024towards , TopV yang2025topv ,Voco-llama ye2024voco , S2-MAD zeng2025s , ToCa zou2025accelerating Vision & Language TransformerLLM & Multimodal LLMEfficiency
+Year2020 - 2021Power-bert goyal2020power , DynamicViT rao2021dynamicvit ,TokenLearner ryoo2021tokenlearner , SpAtten wang2021spatten 2022LTP kim2022learned , SPViT kong2022spvit , EViT liang2022not ,PatchSlim tang2022patch , STTS wang2022efficient , PPT ma2022ppt 2023DCP anagnostidis2023dynamic , ToMe bolya2022token , ToMeSD bolya2023token , PuMer cao2023pumer ,HeatViT dong2023heatvit , Llmlingua jiang2023llmlingua ,DTP nawrot2022efficient , TPS wei2023joint 2024Honeybee cha2024honeybee , FastV chen2024image , TokenPacker li2024tokenpacker ,BRAVE kar2024brave , LLaMA-VID li2024llama , DyDiT zhao2024dynamic - 2025.07DivPrune alvar2025divprune , SeTok wu2024towards , TopV yang2025topv , SparseVLM zhang2024sparsevlm Voco-llama ye2024voco , S2-MAD zeng2025s , ToCa zou2025accelerating ,TALE han2024token Vision & Language TransformerLLM & Multimodal LLMEfficiency
 $\bm{\checkmark}$
-$\bm{\Rightarrow}$  Future
+$\bm{\Rightarrow}$  Future ?
 Figure 1: Timeline of notable method developments for token reduction methods with modality shifts (Vision & Language $\to$ Multimodal LLMs).
-All these strategies aim to speed up inference with neglected performance drops.
+All these strategies aim to speed up inference with negligible performance drops.
 Conversely, we ask: What is the next token reduction paradigm in generative model design that goes beyond test-time accelerations?
@@
 Unlike token reduction in image classification, video compression focuses more on the temporal redundancy within videos, and algorithms are developed to reduce the number of tokens with less computational overhead.
-Various token reduction methods have been investigated for different tasks including video understanding ryoo2021tokenlearner ; wang2022efficient , video editing li2024vidtome , video-text retrieval liu2022ts2 ; shen2024tempme , video action detection chen2023efficient and so on.
+Various token reduction methods have been investigated for different tasks, including video understanding ryoo2021tokenlearner ; sun2025llavascissortokencompressionsemantic ; wang2022efficient , video editing li2024vidtome , video-text retrieval liu2022ts2 ; shen2024tempme , video action detection chen2023efficient , and so on.
 Specifically, STTS wang2022efficient introduces a lightweight framework that dynamically selects the most informative spatial-temporal tokens in video transformers.
@@
 5.2 Token Reduction for Reinforcement Learning
-Reinforcement learning (RL)-driven token reduction technology demonstrates significant potential in multimodal large models, with its core lying in balancing computational efficiency and reasoning accuracy through dynamic reward mechanisms and sparsity constraints.
+Reinforcement learning (RL)-driven token reduction technology demonstrates significant potential in multimodal large models, with its core lying in balancing computational efficiency and reasoning accuracy through dynamic reward mechanisms and sparsity constraints wang2025beyond .
 Under the “Fast-Slow Thinking” framework xiao2025fast , RL can guide hierarchical screening of high-value tokens: The fast path employs sparsity strategies (e.g., rule-based rewards or information density scoring) to filter redundant visual or semantic features, while the slow path focuses on refined reasoning.
 This hierarchical mechanism not only reduces computational overhead by up to 30-50%, but also continuously optimizes token selection strategies via RL’s online feedback.
-For instance, integrating self-play mechanisms with the GRPO algorithm (Group Relative Policy Optimization) enables the generation of diverse candidate paths and selects optimal outputs.
-Furthermore, the sparsity reinforcement paradigm proposed in ZipR1 chen2025zipr1reinforcingtokensparsity highlights RL’s unique advantage in cross-modal generalization.
+The sparsity reinforcement paradigm proposed in ZipR1 chen2025zipr1reinforcingtokensparsity highlights RL’s unique advantage in cross-modal generalization.
 By designing sparsity rewards (e.g., token quantity limits or step compression incentives), models can autonomously generate concise intermediate reasoning chains and even distill such capabilities into lightweight models.
-For example, in visual question answering tasks, models can dynamically extract key image region features while enhancing answer reliability through multi-candidate contrast mechanisms (e.g., parallel samplings), achieving the goal of “low computational cost, high task accuracy.” Looking ahead, this technology could advance toward adaptive sparsity and cross-modal alignment, providing efficient inference foundations for edge computing and real-time interaction scenarios.
-Simultaneously, it will drive MLLMs toward co-evolution of lightweight design and strong robustness.
-5.3 From Prompt Tuning to In-context-learning
+For example, in visual question answering tasks, models can dynamically extract key image region features while enhancing answer reliability through multi-candidate contrast mechanisms (e.g., parallel samplings), achieving the goal of “low computational cost, high task accuracy.”
+Recent findings further reveal that optimizing only a small subset of high-entropy "forking tokens" can match or even surpass full-gradient updates in reasoning tasks wang2025beyond , suggesting that token-level entropy patterns can guide efficient policy updates.
+This entropy-aware strategy reinforces the broader view that sparse, selective token refinement is sufficient for preserving reasoning performance in large models.
+Looking ahead, the integration of such approaches could enhance cross-modal alignment and inference efficiency in real-time and resource-constrained scenarios, supporting a new generation of lightweight yet capable multimodal language agents.
+5.3 From Prompt Tuning to Chain of Thought Reasoning
 Current token reduction efforts for prompts have primarily aimed at compressing prompts for efficiency, often with impressive results jiang2023llmlingua ; mu2023learning .
 Looking forward, token reduction should evolve into enhancing reasoning and maximizing utility per token in context.
-Instead of focusing solely on making prompts shorter, future research should explore how each remaining token can carry more information or trigger more complex inference during in-context learning.
+Instead of focusing solely on making prompts shorter, future research should explore how each remaining token can carry more information or trigger more complex inference during in-context learning and chain of thought reasoning.
 One direction is to alter the generation paradigm itself, for example, training language models to predict multiple tokens per step gloeckle2024better .
-Another idea is to enable deeper internal reasoning without increasing prompt length.
-CoD adams2023sparse introduces an iterative approach using token reduction within fixed-length constraints: Each step incorporates new salient entities into the summary by compressing/abstracting existing content, forcing the model to enhance density without expanding length.
+Another idea is to enable deeper internal reasoning without increasing prompt length adams2023sparse .
+As mentioned in Sec.
+4, long CoT chains can become verbose: excessive reasoning steps may introduce errors or obscure logical clarity, particularly in LLM agents where internal reasoning alternates with external tool use gao2025txagent ; wang2024toolgen .
+Token reduction may serve a critical role in this context by compressing intermediate reasoning into a compact representation.
+For example, approaches based on next-token prediction mu2023learning ; zhang2025lightthinker can distill intermediate thinking chains into a set of dense, information-rich tokens.
+These compressed representations can then replace the full intermediate context and serve as inputs for subsequent reasoning steps.
+This compressed-thinking strategy has two main benefits: 1) reducing error accumulation and keeping the logic clear by focusing on key information, and 2) allowing more reasoning rounds to fit within a fixed context window, enabling deeper multi-step inference without exceeding length limits.
 In summary, the next phase of token reduction research should shift focus from simple prompt compression to reasoning-centric compression.
 Rather than just trimming prompts, we should ask: How can we make each token in the prompt or context do more work for us?
-This involves training models with objectives that reward higher-level inference per token, developing architectures that recycle tokens for multi-step thinking, or dynamically selecting the most salient tokens to keep at each stage of dialogue.
+This involves training models with objectives that reward higher-level inference per token, developing architectures that recycle tokens for multi-step thinking, or dynamically selecting the most salient tokens to keep at each step of reasoning.
 5.4 Complementary to Other Methods
@@
 This process drastically reduces effective vocabulary size, addressing the scalability challenge of 600,000+ medical codes by collapsing redundant representations while preserving critical clinical context.
-Inspired by adaptive tokenization methods for vision duggal2024adaptive ; yan2024elastictok ; kang2023tictok , future EHR tokenization would be adaptive, enabling dynamic representation of patients’ medical histories, where the length of the token series of each patient history would be directly correlated to length and complexity.
+Inspired by adaptive tokenization methods for vision duggal2024adaptive ; kang2023tictok ; yan2024elastictok , future EHR tokenization would be adaptive, enabling the dynamic representation of patients’ medical histories, where the length of the token series for each patient’s history would be directly correlated with the length and complexity.
 Such adaptive tokenization can significantly improve training and inference efficiency across diverse healthcare systems.

### v3

@@
 Despite these advances, this paper argues that token reduction should transcend its traditional efficiency-oriented role in the era of large generative models.
-Instead, we position it as a fundamental principle in generative modeling, critically influencing both model architecture and broader applications.
-Specifically, we contend that across vision, language, and multimodal systems, token reduction can: (i) facilitate deeper multimodal integration and alignment, (ii) mitigate "overthinking" and hallucinations, (iii) maintain coherence over long inputs, and (iv) enhance training stability, etc.
+In this paper, we characterize this mechanism as a fundamental principle in generative modeling, critically influencing both model architecture and broader applications.
+We analyze how token reduction addresses critical challenges in current systems across vision, language, and multimodal, demonstrating its ability to: (i) facilitate deeper multimodal integration and alignment, (ii) mitigate "overthinking" and hallucinations, (iii) maintain coherence over long inputs, and (iv) enhance training stability, etc.
 We reframe token reduction as more than an efficiency measure.
-By doing so, we outline promising future directions, including algorithm design, reinforcement learning-guided token reduction, token optimization for in-context learning, and broader ML and scientific domains.
-We highlight its potential to drive new model architectures and learning strategies that improve robustness, increase interpretability, and better align with the objectives of generative modeling.
+By doing so, we outline promising future directions, including algorithm design, reinforcement learning-guided token reduction, token optimization for in-context learning, agentic framework design, and broader ML and scientific domains.
 [^1] We collected a list of token reduction papers at: Awesome-Collection-Token-Reduction.
 1 Introduction
-Transformer-based generative models brown2020language ; devlin2019bert ; NEURIPS2021_854d9fca ; vaswani2017attention have emerged as dominant deep learning architectures across vision, language, and multimodal tasks, due to their ability to process long sequences of tokens, which are the fundamental representational units derived from raw data such as subwords in language or image patches in vision.
+Transformer-based generative models Brown et al.
+(2020); Devlin et al.
+(2019); Han et al.
+(2021); Vaswani et al.
+(2017) have emerged as dominant deep learning architectures across vision, language, and multimodal tasks, due to their ability to process long sequences of tokens, which are the fundamental representational units derived from raw data such as subwords in language or image patches in vision.
 As these models are applied to increasingly complex real-world tasks, the input sequence lengths of both the models and their training datasets continue to grow.
@@
 Token reduction addresses this challenge by reducing the number of tokens processed during inference.
-By pruning or merging tokens, token reduction decreases computational cost and accelerates runtime, offering a practical solution for improving the efficiency of generative models.
+By pruning or merging tokens, token reduction Guo et al.
+(2025); Han et al.
+(2026); Huang et al.
+(2024a); Hyun et al.
+(2025); Jiang et al.
+(2025b); Kim et al.
+(2025); Li et al.
+(2025a); Liu et al.
+(2025b); Wu et al.
+(2025e); Zhang et al.
+(2025d) reduces computational cost and accelerates runtime, providing a practical solution for enhancing generative efficiency Kim et al.
+(2025); Liu et al.
+(2025c); Tang et al.
+(2025); Wu et al.
+(2025a); Xing et al.
+(2025); Yao et al.
+(2025); Zhang et al.
+(2025a); Zhang and Fu (2025); Zhang et al.
+(2025e); Zhuang et al.
+(2025).
 Token reduction has been widely adopted in computer vision, language processing, and multimodal tasks.
-In vision transformers, it has primarily been used to reduce computational cost by removing visually redundant tokens bolya2022token ; kong2022spvit ; liang2022not ; rao2021dynamicvit .
-While effective in decreasing token-level computations and accelerating inference, this approach can discard subtle but important visual features, especially in dense prediction tasks such as segmentation and object detection, leading to significant performance degradation.
-In language models, token reduction has commonly been implemented through early-exit mechanisms and token-skipping strategies lin2025boosting ; wu2024accelerating , which reduce the number of intermediate tokens processed and thus lower computational overhead.
-Similarly, multimodal large language models (MLLMs) apply visual token pruning primarily during the prefill stage chen2024image , where adaptive attention patterns are learned in the early layers to prune tokens in later stages.
-Despite progress in existing work, token reduction is still predominantly viewed as a means of improving computational efficiency, primarily by reducing the number of tokens to minimize associated computations and accelerate inference.
-Such an efficiency-only mindset has critical limitations.
-Naive pruning methods may discard informative tokens, thereby degrading model understanding and performance liang2022not ; zhan2024exploring ; zhan-etal-2024-rethinking-token .
-Furthermore, token reduction is commonly treated as a post hoc optimization, rather than being integrated into the core design and training of the model chen2024image .
+In vision transformers, it has primarily been used to reduce computational cost by removing visually redundant tokens Bergner et al.
+(2025); Bolya et al.
+(2023); Fang et al.
+(2025a); Hong and Liu (2025); Kong et al.
+(2022); Lei et al.
+(2025); Liang et al.
+(2022b); Rao et al.
+(2021).
+In language models, token reduction has commonly been implemented through early-exit mechanisms and token-skipping strategies Lin et al.
+(2025); Wu et al.
+(2024), which reduce the number of intermediate tokens processed and thus lower computational overhead.
+Similarly, multimodal large language models (MLLMs) apply visual token pruning primarily during the prefill stage Chen et al.
+(2024), where adaptive attention patterns are learned in the early layers to prune tokens in later stages.
+Despite progress, token reduction is predominantly viewed as a post-hoc efficiency optimization Liu et al.
+(2025d); Shao et al.
+(2025), primarily by reducing the number of tokens to minimize associated computations and accelerate inference.
+Such an efficiency-only mindset has limitations.
+Naive pruning methods may discard informative tokens, thereby degrading model understanding and performance Liang et al.
+(2022b); Zhan et al.
+(2024a, b).
+Furthermore, token reduction is commonly treated as a post hoc optimization, rather than being integrated into the core design and training of the model Chen et al.
+(2024).
 In this paper, we argue that viewing token reduction purely from an efficiency perspective is fundamentally limited.
@@
 (iv) Noisy or irrelevant tokens introduced during training slow down convergence and harm model stability.
-By learning to intelligently select, merge, or compress tokens based on their contribution to generation objectives-rather than solely on raw redundancy-models can simultaneously reduce computational load, improve robustness, and enhance interpretability and alignment.
+By learning to intelligently select, merge, or compress tokens based on their contribution to generation objectives, rather than solely on raw redundancy, models can simultaneously reduce computational load, improve robustness, and enhance interpretability and alignment.
 This paper makes the following three key contributions:
-  • We provide a comprehensive analysis of the limitations inherent in current efficiency-centric token pruning methods across vision, language, and multimodal generative models, highlighting critical shortcomings in these widely adopted strategies.
+  • We categorize existing token reduction methods by their functional objective, identifying a transition from efficiency-centric optimizations to task-aware enhancements in vision, language, and multimodal domains.
   • We identify core challenges faced by modern generative models including insufficient visual representation, semantic misalignment, overthinking in reasoning, and training instability.
 We then demonstrate how principled token reduction strategies can effectively mitigate these issues.
-  • We propose a roadmap for future research on token reduction, including directions for method design, reinforcement learning-guided token selection, adaptive in-context compression, and hardware-algorithm co-design, etc.
+  • We outline a roadmap for future research on token reduction, including directions for method design, reinforcement learning-guided token selection, adaptive in-context compression, and hardware-algorithm co-design, etc.
 These directions aim to support the development of next-generation generative architectures that are both robust and efficient.
@@
 Classification serves as a fundamental task for vision models and token reduction techniques have been widely applied in it due to its simplicity and versatility.
-It has been widely explored from various aspects bolya2022token ; liang2022not ; rao2021dynamicvit ; wu2023ppt ; zeng2024m2m ; kong2023peeling ; zeng2025token .
-Specifically, DynamicViT rao2021dynamicvit devises a lightweight module to predict the importance score of each token, thereby pruning unimportant tokens.
-SPViT kong2022spvit introduces a soft pruning technique, which integrates the less informative tokens generated by the selector module into a package token that will participate in subsequent calculations rather than being completely discarded.
-EViT liang2022not identifies attentive tokens from the attention map, enabling token pruning without additional parameters.
-ToMe bolya2022token merges tokens with similarity based on bipartite matching to maintain information utility.
-PPT wu2023ppt analyzes the statistic data between layers and adaptively employs token pruning and merging within layers to achieve higher acceleration performance.
-Year2020 - 2021Power-bert goyal2020power , DynamicViT rao2021dynamicvit ,TokenLearner ryoo2021tokenlearner , SpAtten wang2021spatten 2022LTP kim2022learned , SPViT kong2022spvit , EViT liang2022not ,PatchSlim tang2022patch , STTS wang2022efficient , PPT ma2022ppt 2023DCP anagnostidis2023dynamic , ToMe bolya2022token , ToMeSD bolya2023token , PuMer cao2023pumer ,HeatViT dong2023heatvit , Llmlingua jiang2023llmlingua ,DTP nawrot2022efficient , TPS wei2023joint 2024Honeybee cha2024honeybee , FastV chen2024image , TokenPacker li2024tokenpacker ,BRAVE kar2024brave , LLaMA-VID li2024llama , DyDiT zhao2024dynamic - 2025.07DivPrune alvar2025divprune , SeTok wu2024towards , TopV yang2025topv , SparseVLM zhang2024sparsevlm Voco-llama ye2024voco , S2-MAD zeng2025s , ToCa zou2025accelerating ,TALE han2024token Vision & Language TransformerLLM & Multimodal LLMEfficiency
-$\bm{\checkmark}$
-$\bm{\Rightarrow}$  Future ?
-Figure 1: Timeline of notable method developments for token reduction methods with modality shifts (Vision & Language $\to$ Multimodal LLMs).
+It has been widely explored from various aspects Bolya et al.
+(2023); Liang et al.
+(2022b); Rao et al.
+(2021); Wu et al.
+(2023); Zeng and Yu (2024); Kong et al.
+(2023); Zeng et al.
+(2025a).
+Specifically, DynamicViT Rao et al.
+(2021) devises a lightweight module to predict the importance score of each token, thereby pruning unimportant tokens.
+SPViT Kong et al.
+(2022) introduces a soft pruning technique, which integrates the less informative tokens generated by the selector module into a package token that will participate in subsequent calculations rather than being completely discarded.
+EViT Liang et al.
+(2022b) identifies attentive tokens from the attention map, enabling token pruning without additional parameters.
+ToMe Bolya et al.
+(2023) merges tokens with similarity based on bipartite matching to maintain information utility.
+PPT Wu et al.
+(2023) analyzes the statistic data between layers and adaptively employs token pruning and merging within layers to achieve higher acceleration performance.
+Year2020 - 2021Power-bert Goyal et al.
+(2020), DynamicViT Rao et al.
+(2021),TokenLearner Ryoo et al.
+(2021), SpAtten Wang et al.
+(2021)2022LTP Kim et al.
+(2022), SPViT Kong et al.
+(2022), EViT Liang et al.
+(2022b),PatchSlim Tang et al.
+(2022), STTS Wang et al.
+(2022), PPT Ma et al.
+(2022)2023DCP Anagnostidis et al.
+(2023), ToMe Bolya et al.
+(2023), ToMeSD Bolya and Hoffman (2023), PuMer Cao et al.
+(2023),HeatViT Dong et al.
+(2023), Llmlingua Jiang et al.
+(2023),DTP Nawrot et al.
+(2023), TPS Wei et al.
+(2023)2024Honeybee Cha et al.
+(2024), FastV Chen et al.
+(2024), LLaMA-VID Li et al.
+(2024c), TokenPacker Li et al.
+(2024a), BRAVE Kar et al.
+(2024), DyDiT Zhao et al.
+(2025), TALE Han et al.
+(2025)2025DivPrune Alvar et al.
+(2025), SeTok Wu et al.
+(2025b), SparseVLM Zhang et al.
+(2025g) TopV Yang et al.
+(2025a), Voco-llama Ye et al.
+(2025), FastDriveVLA Cao et al.
+(2025),EfficientVLA Yang et al.
+(2025c), S2-MAD Zeng et al.
+(2025b), LatentMAS Zou et al.
+(2025b)Vision & Language TransformerLLM & Multimodal LLM & AgentsEfficiency
+[…]
+[…]
+Figure 1: Timeline of notable method developments for token reduction methods with modality shifts (Vision & Language $\to$ Multimodal LLMs & Agents).
 All these strategies aim to speed up inference with negligible performance drops.
@@
 Unlike token reduction in image classification, video compression focuses more on the temporal redundancy within videos, and algorithms are developed to reduce the number of tokens with less computational overhead.
-Various token reduction methods have been investigated for different tasks, including video understanding ryoo2021tokenlearner ; sun2025llavascissortokencompressionsemantic ; wang2022efficient , video editing li2024vidtome , video-text retrieval liu2022ts2 ; shen2024tempme , video action detection chen2023efficient , and so on.
-Specifically, STTS wang2022efficient introduces a lightweight framework that dynamically selects the most informative spatial-temporal tokens in video transformers.
-Tokenlearner ryoo2021tokenlearner proposes an adaptive tokenization module that learns a handful of informative spatial-temporal tokens, significantly reducing computational costs.
-EVAD chen2023efficient selectively drops irrelevant spatial-temporal tokens in non-keyframes while preserving keyframe and motion-relevant tokens, and then refines actor features using a context-aware decoder to maintain accuracy with reduced computations.
+Various token reduction methods have been investigated for different tasks, including video understanding Ryoo et al.
+(2021); Sun et al.
+(2025); Wang et al.
+(2022), video editing Li et al.
+(2024b), video-text retrieval Liu et al.
+(2022); Shen et al.
+(2025a), video action detection Chen et al.
+(2023), and so on.
+Specifically, STTS Wang et al.
+(2022) introduces a lightweight framework that dynamically selects the most informative spatial-temporal tokens in video transformers.
+Tokenlearner Ryoo et al.
+(2021) proposes an adaptive tokenization module that learns a handful of informative spatial-temporal tokens, significantly reducing computational costs.
+EVAD Chen et al.
+(2023) selectively drops irrelevant spatial-temporal tokens in non-keyframes while preserving keyframe and motion-relevant tokens, and then refines actor features using a context-aware decoder to maintain accuracy with reduced computations.
 Generative Tasks.
-Token reduction in generative tasks ju2024turbo aims to accelerate generative models through the efficient utilization of tokens.
-It can be applied to both diffusion models bolya2023token ; wang2024attention and diffusion transformers zou2025accelerating .
-Specifically, ToMeSD bolya2023token exploits natural redundancy in generated images by merging redundant tokens, successfully extending token merging to stable diffusion with simple unmerging.
-DyDiT zhao2024dynamic reduces redundancy with a Timestep-wise Dynamic Width approach to adopt model width conditioned on the generation timesteps, and a Spatial-wise Dynamic Token strategy to avoid redundant computations at unnecessary spatial locations.
+Token reduction in generative tasks Ju et al.
+(2024) aims to accelerate generative models through the efficient utilization of tokens.
+It can be applied to both diffusion models Bolya and Hoffman (2023); Lu et al.
+(2025); Wang et al.
+(2024a) and diffusion transformers Gao et al.
+(2023); Zou et al.
+(2025a).
+Specifically, ToMeSD Bolya and Hoffman (2023) exploits natural redundancy in generated images by merging redundant tokens, successfully extending token merging to stable diffusion with simple unmerging.
+DyDiT Zhao et al.
+(2025) reduces redundancy with a Timestep-wise Dynamic Width approach to adopt model width conditioned on the generation timesteps, and a Spatial-wise Dynamic Token strategy to avoid redundant computations at unnecessary spatial locations.
 2.2 Token Reduction in Language Models
-Token reduction strategies in language modeling have evolved from early optimizations for BERT huang2022pyramid ; kim2020length ; kim2022learned ; kim2023leap ; ye2021tr to techniques specifically designed for LLMs.
-PoWER-BERT goyal2020power introduces progressive word-vector elimination by removing redundant token representations based on self-attention dynamics, improving inference efficiency.
-Learned token pruning kim2022learned extends this approach by learning attention-based thresholds to adaptively prune uninformative tokens, thereby reducing computational costs while preserving model performance.
+Token reduction strategies in language modeling have evolved from early optimizations for BERT Huang et al.
+(2022); Kim and Cho (2021); Kim et al.
+(2022, 2023); Ye et al.
+(2021) to techniques specifically designed for LLMs.
+PoWER-BERT Goyal et al.
+(2020) introduces progressive word-vector elimination by removing redundant token representations based on self-attention dynamics, improving inference efficiency.
+Learned token pruning Kim et al.
+(2022) extends this approach by learning attention-based thresholds to adaptively prune uninformative tokens, thereby reducing computational costs while preserving model performance.
 In LLMs, token reduction must account for the constraints of autoregressive decoding across diverse downstream tasks.
-Dynamic pooling methods anagnostidis2023dynamic ; tao2025saliency adjust token representations on the fly during inference to reduce redundancy.
-Prompt compression techniques fu2024lazyllm ; ge2023context ; jiang2023llmlingua aim to reduce computational overhead by compressing the input prompt before generation.
-Selective decoding approaches fu2024lazyllm ; wingate2022prompt reduce per-step inference costs by computing key-value pairs only for tokens critical to predicting the next token.
-In multi-agent systems, S2-MAD zeng2025s proposes a sparsification mechanism that limits unnecessary token exchanges between agents, reducing communication costs and improving the efficiency of collaborative reasoning.
+Dynamic pooling methods Anagnostidis et al.
+(2023); Tao et al.
+(2025) adjust token representations on the fly during inference to reduce redundancy.
+Prompt compression techniques Fu et al.
+(2024); Ge et al.
+(2024); Jiang et al.
+(2023) aim to reduce computational overhead by compressing the input prompt before generation.
+Selective decoding approaches Fu et al.
+(2024); Wingate et al.
+(2022) reduce per-step inference costs by computing key-value pairs only for tokens critical to predicting the next token.
+In multi-agent systems, S2-MAD Zeng et al.
+(2025b) proposes a sparsification mechanism that limits unnecessary token exchanges between agents, reducing communication costs and improving the efficiency of collaborative reasoning.
 2.3 Token Reduction in Multimodal LLMs
-Recent work has explored visual token pruning by addressing attention inefficiencies of deep transformer layers in MLLMs alvar2025divprune ; arif2025hired ; lin2025boosting ; shang2024llava .
-Specifically, FastV chen2024image shows that deeper vision-language layers expend significant computations on redundant image tokens.
+Recent work has explored visual token pruning by addressing attention inefficiencies of deep transformer layers in MLLMs Alvar et al.
+(2025); Arif et al.
+(2025); Lin et al.
+(2025); Shang et al.
+(2024).
+Specifically, FastV Chen et al.
+(2024) shows that deeper vision-language layers expend significant computations on redundant image tokens.
 To address this, a lightweight module is adopted to adaptively prune these tokens, reducing inference overheads in subsequent stages.
-A complementary approach modifies vision feature extractors or projectors to output a smaller set of highly informative image tokens, effectively distilling the input into a compressed representation cai2024matryoshka ; cha2024honeybee ; kar2024brave ; li2024tokenpacker ; li2024mini ; ye2024voco .
+A complementary approach modifies vision feature extractors or projectors to output a smaller set of highly informative image tokens, effectively distilling the input into a compressed representation Cai et al.
+(2025); Cha et al.
+(2024); Kar et al.
+(2024); Li et al.
+(2024a, d); Ye et al.
+(2025).
 However, efficiency gains from these prefill stages often fade during the decoding phase, where per-token computations dominate.
-To overcome this, recent methods jointly optimize token reduction during both prefill and decoding stages, ensuring sustained speedups throughout inference huang2024dynamic ; song2024less .
+To overcome this, recent methods jointly optimize token reduction during both prefill and decoding stages, ensuring sustained speedups throughout inference Huang et al.
+(2025); Song et al.
+(2025).
+Furthermore, the scope of token reduction has recently expanded to Vision-Language-Action (VLA) models, optimizing efficiency for real-time robotic manipulation and autonomous driving Cao et al.
+(2025); Jiang et al.
+(2025c); Pei et al.
+(2025); Yang et al.
+(2025c).
 In Fig.
-1, we present a timeline of notable developments in token reduction methods, illustrating the shift from early applications in ViT and BERT-based models to more recent advances in LLMs and MLLMs.
+1, we present a timeline of notable developments in token reduction methods, illustrating the shift from early applications in ViT and BERT-based models to more recent advances in LLMs, MLLMs and Agent systems.
 3 Problem Formulation
-In modern generative models Qwen2.5VL ; 2024llama3 ; liu2023llava ; openai2024gpt4 ; Peebles2022DiT , a token denotes one fundamental unit of input or representation, typically encoded as a vector.
+In modern generative models Bai et al.
+(2025); Grattafiori et al.
+(2024); Liu et al.
+(2023); OpenAI et al.
+(2024); Peebles and Xie (2023), a token denotes one fundamental unit of input or representation, typically encoded as a vector.
 For example, a token might correspond to a subword in language, a patch in an image, or an embedding of a time step in audio.
@@
 Token reduction refers to any operation that compresses the token sequence to $M$ tokens (with $M<N$) by removing or consolidating tokens while aiming to preserve the original information.
-Broadly, token reduction methods fall into the following categories: 1) Token pruning methods bolya2022token that remove entire unimportant tokens, simply dropping them from the sequence; 2) Token merging methods bolya2022token ; bolya2023token which fuse information from multiple tokens into fewer tokens, effectively compressing the sequence by merging similar or related tokens; 3) Hybrid strategies cao2023pumer ; kim2024token ; wu2023ppt that combine pruning and merging within a unified framework; 4) Token distillation approaches cai2024matryoshka ; mu2023learning which integrate rich information across longer input sequences or multiple modalities into fewer condensed tokens, enabling efficient cross-modal interactions and long-context reasoning in LLMs and MLLMs.
+Broadly, token reduction methods fall into four categories: 1) Token pruning methods Bolya et al.
+(2023) that remove entire unimportant tokens, simply dropping them from the sequence; 2) Token merging methods Bolya et al.
+(2023); Bolya and Hoffman (2023) which fuse information from multiple tokens into fewer tokens, effectively compressing the sequence by merging similar or related tokens; 3) Hybrid strategies Cao et al.
+(2023); Kim et al.
+(2024); Wu et al.
+(2023) that combine pruning and merging within a unified framework; 4) Token distillation approaches Cai et al.
+(2025); Mu et al.
+(2023) which integrate rich information across longer input sequences or multiple modalities into fewer condensed tokens, enabling efficient cross-modal interactions and long-context reasoning in LLMs and MLLMs.
 A core challenge in token reduction is the determination of tokens to be pruned or merged.
-There are various importance criteria and scoring mechanisms to rank token significance, including attention-based heuristics liang2022not , gradient or loss-based criteria huang2024dynamic , clustering haurum2024agglomerative , and learned predictors rao2021dynamicvit .
+There are various importance criteria and scoring mechanisms to rank token significance, including attention-based heuristics Liang et al.
+(2022b), gradient or loss-based criteria Huang et al.
+(2025), clustering Haurum et al.
+(2024), and learned predictors Rao et al.
+(2021).
 From a purely efficiency-oriented perspective, token reduction delivers substantial computational efficiency gains by reducing the quadratic computation cost from $\mathcal{O}(N^{2})$ to $\mathcal{O}(M^{2})$ in attention mechanisms.
 By eliminating redundant tokens and processing fewer computations during inference, it effectively accelerates the inference speed and improves the model throughput, which is crucial for latency-sensitive tasks or real-time applications.
-Furthermore, it also shrinks the memory footprint for activations and gradients (e.g., key/value caches), alleviating memory usage for both inference and training, which is particularly beneficial for wide deployments on resource-limited platforms.
+Furthermore, it reduces the memory footprint of activations and gradients (e.g., key/value caches), alleviating memory usage for both inference and training, which is particularly beneficial for wide-scale deployments on resource-limited platforms.
+We present more theoretical detail in the Appendix A
 However, as stated in this position paper, token reduction can benefit models in multiple ways beyond efficiency, which will be introduced in detail in the following sections.
-4 Challenges and Positions
-In this section, we establish token reduction as a foundational mechanism for addressing critical challenges in modern generative systems.
-We claim five core challenges across modalities: visual representation sparsity, semantic misalignment, reasoning redundancy, training instability, and long-context overload.
+4 Core Roles and Challenges
+In this section, we discuss token reduction as a foundational mechanism for addressing critical challenges in modern generative systems.
+We categorize five core challenges across modalities: visual representation sparsity, semantic misalignment, reasoning redundancy, training instability, and long-context overload.
 We demonstrate how principled token reduction strategies intrinsically address these issues through dynamic token-semantic co-optimization.
@@
 MLLMs often suffer from noisy visual inputs that impede fine-grained understanding.
-We outline key challenges in MLLM visual reasoning: ① Text-Visual Attention Shift: Due to the rotary positional embeddings in LLM decoders, later text tokens disproportionately attend to spatially lower image regions hong2024token , shifting attention away from semantically important areas (e.g., objects at the top of an image); ② Visual Redundancy: Empirical studies lin2025boosting ; wu2024accelerating show that beyond the first few layers, many image tokens contribute little new information, ③ Task-Guided Focus in VQA: In multimodal question answering, the question itself pinpoints relevant image regions (e.g., "kitten color" directs focus to the kitten patch), implying that many image tokens are unnecessary for correct answers song2024less .
-Therefore, we position token reduction as a representation-learning optimization: selecting the subset of tokens that preserves informative visual representation.
-For example, VisPruner zhang2025textvisualattentionexploitingvisual identifies high-value tokens using visual-encoder attention and removes duplicates via clustering to ensure diversity.
-VTW lin2025boosting observes that visual information migrates into text tokens within early layers; it therefore withdraws all visual tokens after a chosen layer based on KL-divergence criteria.
-TRIM song2024less leverages the CLIP metric and IQR scoring function to adaptively select image tokens that are crucial for answering questions, while an aggregated token is used to retain additional image information.
+We outline key challenges in MLLM visual reasoning: ① Text-Visual Attention Shift: Due to the rotary positional embeddings in LLM decoders, later text tokens disproportionately attend to spatially lower image regions Hong et al.
+(2024), shifting attention away from semantically important areas (e.g., objects at the top of an image); ② Visual Redundancy: Empirical studies Lin et al.
+(2025); Wu et al.
+(2024) show that beyond the first few layers, many image tokens contribute little new information, ③ Task-Guided Focus in VQA: In multimodal question answering, the question itself pinpoints relevant image regions (e.g., "kitten color" directs focus to the kitten patch), implying that many image tokens are unnecessary for correct answers Song et al.
+(2025).
+Therefore, token reduction can serve as a representation-learning optimization: selecting the subset of tokens that preserves informative visual representation.
+For example, VisPruner Zhang et al.
+(2025c) identifies high-value tokens using visual-encoder attention and removes duplicates via clustering to ensure diversity.
+VTW Lin et al.
+(2025) observes that visual information migrates into text tokens within early layers; it therefore withdraws all visual tokens after a chosen layer based on KL-divergence criteria.
+TRIM Song et al.
+(2025) leverages the CLIP metric and IQR scoring function to adaptively select image tokens that are crucial for answering questions, while an aggregated token is used to retain additional image information.
 4.2 Better Multimodal Token Alignment
@@
 Token reduction offers a promising solution by selecting visual tokens based on semantic importance, thereby producing a compact set of tokens that better align with language representations.
-Specifically, SeTok wu2024towards dynamically clusters visual features into semantically meaningful tokens using a density-peak algorithm, which determines both the number and structure of token groupings per image.
+Specifically, SeTok Wu et al.
+(2025b) dynamically clusters visual features into semantically meaningful tokens using a density-peak algorithm, which determines both the number and structure of token groupings per image.
 This approach preserves both high- and low-frequency semantics, substantially improving concept-level alignment and downstream task performance.
-M3 cai2024matryoshka introduces a hierarchical token structure that captures coarse-to-fine semantic granularity, allowing different levels of abstraction to be selectively retained depending on task needs.
+M3 Cai et al.
+(2025) introduces a hierarchical token structure that captures coarse-to-fine semantic granularity, allowing different levels of abstraction to be selectively retained depending on task needs.
 4.3 Reduce Overthinking in Reasoning
@@
 In the context of language models, overthinking refers to generating excessively long or convoluted chains of reasoning that go beyond what is necessary to reach a correct answer.
-An LLM may produce verbose, repetitive, or even self-contradictory explanations when it fails to converge on a solution-often due to uncertainty sui2025stop ; wang2025harnessing .
-Such extended reasoning trajectories are inefficient and recent studies show that state-of-the-art reasoners can consume over 15,000 tokens to solve math problems that could be addressed with a concise chain-of-thought (CoT) of just a few hundred tokens hou2025thinkprune .
-Mitigating overthinking is thus crucial for improving both accuracy and efficiency in generative modeling.
-By reducing unnecessary tokens during reasoning, LLMs can focus on salient steps, aligning generation with a more concise and logical trajectory.
-CoT-Influx huang2023fewer introduces a CoT pruning strategy in which concise reasoning examples are included in the prompt.
+An LLM may produce verbose, repetitive, or even self-contradictory explanations when it fails to converge on a solution-often due to uncertainty Sui et al.
+(2025); Wang et al.
+(2025a).
+Such extended reasoning trajectories are inefficient and recent studies show that state-of-the-art reasoners can consume over 15,000 tokens to solve math problems that could be addressed with a concise chain-of-thought (CoT) of just a few hundred tokens Hou et al.
+(2025).
+This issue is particularly acute in LLM agents, where internal reasoning alternates with external tool use Gao et al.
+(2025a); Wang et al.
+(2024b); excessive steps can obscure logical clarity and lead to error accumulation.
+Mitigating overthinking is thus crucial.
+By trimming unnecessary tokens, LLMs can focus on salient steps, aligning generation with a more concise trajectory.
+CoT-Influx Huang et al.
+(2024b) introduces a CoT pruning strategy in which concise reasoning examples are included in the prompt.
 By pruning unimportant tokens from these examples, more reasoning demonstrations can fit into the context window, surprisingly leading to improved math reasoning accuracy.
-TokenSkip xia2025tokenskip enables LLMs to skip less important tokens within CoT sequences and learn shortcuts between critical reasoning steps.
+TokenSkip Xia et al.
+(2025) enables LLMs to skip less important tokens within CoT sequences and learn shortcuts between critical reasoning steps.
 This allows for controllable CoT compression with adjustable compression ratios, enabling models to automatically trim redundant tokens during reasoning.
@@
 MLLMs, which reason over text and other modalities, face similar overthinking issues.
-In vision-language tasks, overthinking often manifests as excessive processing of visual tokens or overly detailed image descriptions, resulting in inefficiency and potential confusion chen2025zipr1reinforcingtokensparsity .
+In vision-language tasks, overthinking often manifests as excessive processing of visual tokens or overly detailed image descriptions, resulting in inefficiency and potential confusion Chen et al.
+(2025).
 Token reduction techniques in MLLMs aim to promote more focused and sparse reasoning over multimodal inputs.
-For example, FAST xiao2025fast rewards shorter-than-average token sequences for correct answers, while allowing longer reasoning for more complex tasks.
+For example, FAST Xiao et al.
+(2025a) rewards shorter-than-average token sequences for correct answers, while allowing longer reasoning for more complex tasks.
 It also adjusts policy optimization constraints to tighten output exploration for simple tasks (thus reducing unnecessary tokens) and loosen it for harder ones to allow deeper reasoning.
 Together, these strategies reduce overthinking in straightforward cases, boosting efficiency while preserving effective reasoning depth for complex scenarios.
-4.4 Improve Training Stability
-While token reduction has traditionally been employed as a post-training optimization to enhance inference efficiency, recent research indicates its potential to significantly improve training stability when integrated into the pre-training phase gao2023masked ; li2024pruning ; lin2024not , suggesting that selective token utilization during training can lead to more robust and effective model learning.
-One notable approach is Rho-1 lin2024not , which involves scoring tokens based on their alignment with a desired distribution using a reference model and then focusing the training loss on tokens with higher scores.
+4.4 Improve Training Stability & Efficiency
+While token reduction has traditionally been employed as a post-training optimization to enhance inference efficiency, recent research indicates its potential to significantly improve training stability when integrated into the pre-training phase Gao et al.
+(2023); Li et al.
+(2025c); Lin et al.
+(2024), suggesting that selective token utilization during training can lead to more robust model learning.
+One notable approach is Rho-1 Lin et al.
+(2024), which involves scoring tokens based on their alignment with a desired distribution using a reference model and then focusing the training loss on tokens with higher scores.
 Therefore, it effectively filters out noisy or less informative tokens, leading to faster convergence and improved performance.
-UPFT ji2025first emphasizes the importance of initial reasoning steps in training.
+UPFT Ji et al.
+(2025) emphasizes the importance of initial reasoning steps in training.
 By reducing the number of training tokens, UPFT encourages the model to focus on the initial prefix substrings of reasoning trajectories, which are often more stable and contain crucial information.
 This focus helps the model avoid being influenced by subsequent complex or potentially erroneous information, thereby improving training stability.
-However, how to design token reduction algorithms that are tightly integrated with training procedures such as in modern methods like GRPO shao2024deepseekmath , remains underexplored.
+Additionally, integrating token reduction with training procedures like GRPO Shao et al.
+(2024) is gaining traction; for instance, recent work reveals that optimizing only a subset of high-entropy ”forking tokens” matches full-gradient updates Wang et al.
+(2025b), suggesting that entropy patterns can effectively guide efficient policy learning.
 Future research should investigate specialized approaches that incorporate token reduction directly into training objectives, enabling models to learn to prioritize or discard tokens in a task-aware and gradient-aligned manner.
@@
 By doing so, models preserve core information such as key events, central themes, or task-specific facts, while significantly decreasing cognitive load.
-For example, AutoCompressors chevalier2023adapting trains pre-trained LLMs to compress long contexts into compact summary tokens, reducing token length by orders of magnitude to extend context windows and speed up inference.
-TokenSwift wu2025hours reduces the effective number of tokens that the model dynamically processes during generation by using multi-token parallel generation and n-gram retrieval for token reutilization, therefore enabling efficient ultra-long sequence generation (up to 100K tokens).
+For example, AutoCompressors Chevalier et al.
+(2023) trains pre-trained LLMs to compress long contexts into compact summary tokens, reducing token length by orders of magnitude to extend context windows and speed up inference.
+TokenSwift Wu et al.
+(2025d) reduces the effective number of tokens that the model dynamically processes during generation by using multi-token parallel generation and n-gram retrieval for token reutilization, therefore enabling efficient ultra-long sequence generation (up to 100K tokens).
 Video-based MLLMs.
@@
 ③ Preserving semantic integrity: it facilitates feasible processing of extremely long sequences in learning while preserving semantic integrity.
-④ Multi-modal alignment: token reduction distills visual information into a compact, semantically aligned form, thereby efficiently bridging the gap between language and vision liu2025hybrid .
+④ Multi-modal alignment: token reduction distills visual information into a compact, semantically aligned form, thereby efficiently bridging the gap between language and vision Liu et al.
+(2025e).
 By doing so, it effectively addresses the challenges posed by the low abstractness and lack of guidance inherent in raw visual inputs, which are the root causes of semantic misalignment and optimization ambiguity in multi-modal models.
-Recent works illustrate these principles: HICom liu2025hybrid conducts conditional token compression at local and global levels using user instructions as guidance to retain instruction-relevant visual information while reducing computational burden.
-Video-XL-Pro liu2025video employs reconstructive token compression with a dynamic token synthesizer and semantic-guided masking to generate compact yet comprehensive video tokens for improved MLLM performance and efficiency.
+Recent works illustrate these principles: HICom Liu et al.
+(2025e) conducts conditional token compression at local and global levels using user instructions as guidance to retain instruction-relevant visual information while reducing computational burden.
+Video-XL-Pro Liu et al.
+(2025a) employs reconstructive token compression with a dynamic token synthesizer and semantic-guided masking to generate compact yet comprehensive video tokens for improved MLLM performance and efficiency.
 5 Future Directions
-In this section, we propose eight promising directions for token reduction beyond the efficiency benefits, organized into three categories: (i) Algorithmic Innovations (Sec.
-5.1$\sim$5.4), (ii) Application Innovations (Sec.
-5.5$\sim$ 5.7), and (iii) Hardware-Algorithm Co-Design (Sec.
-5.8).
+[…]
 5.1 Design of New Algorithms
@@
 It is critical to re-evaluate how token importance is defined and measured.
-More robust and unbiased scoring mechanisms can be developed, such as predictors akhauri2025tokenbutler or meta-learning frameworks that go beyond attention-based proxies.
+More robust and unbiased scoring mechanisms can be developed, such as predictors Akhauri et al.
+(2025) or meta-learning frameworks that go beyond attention-based proxies.
 These models should capture downstream utility with minimal supervision, enabling adaptive pruning across tasks and domains.
 Constructive Token Compression.
-Token reduction can shift from purely eliminative pruning to strategies that merge spatially or semantically similar tokens into compact summary vectors li2024tokenpacker .
+Token reduction can shift from purely eliminative pruning to strategies that merge spatially or semantically similar tokens into compact summary vectors Li et al.
+(2024a).
 Mitigating Position Bias.
-In MLLMs, attention-based pruning methods (e.g., FastV) often rely on attention scores from a fixed query token, leading to retained tokens concentrating in specific image regions (e.g., lower corner) wen2025token with potential position bias.
+In MLLMs, attention-based pruning methods (e.g., FastV) often rely on attention scores from a fixed query token, leading to retained tokens concentrating in specific image regions (e.g., lower corner) Wen et al.
+(2025) with potential position bias.
 Future methods should preserve spatial diversity by enforcing structural uniformity in retained tokens to improve robustness on visual tasks.
@@
 Pruning decisions in MLLMs should be guided by inter-modality dependencies, rather than made independently for each modality.
-For example, text-guided pruning of visual tokens can improve alignment between modalities cao2024madtp .
+For example, text-guided pruning of visual tokens can improve alignment between modalities Cao et al.
+(2024).
 The design should account for joint representations and semantic correspondence across all relevant inputs.
@@
 Token reduction should consider both the prefill stage and decoding phase for LLMs.
-This includes dynamically managing the sparsity of KV caches and selectively updating generated tokens, sustaining efficiency gains throughout the entire inference process  huang2024dynamic .
+This includes dynamically managing the sparsity of KV caches and selectively updating generated tokens, sustaining efficiency gains throughout the entire inference process  Huang et al.
+(2025).
 Hardware-Algorithm Co-Design.
 Token pruning can explore custom hardware and compiler optimizations that take advantage of dynamic token sparsity patterns (e.g., irregular memory access and conditional computation) to maximize throughput and energy efficiency as detailed in Sec.
-5.8.
-5.2 Token Reduction for Reinforcement Learning
-Reinforcement learning (RL)-driven token reduction technology demonstrates significant potential in multimodal large models, with its core lying in balancing computational efficiency and reasoning accuracy through dynamic reward mechanisms and sparsity constraints wang2025beyond .
-Under the “Fast-Slow Thinking” framework xiao2025fast , RL can guide hierarchical screening of high-value tokens: The fast path employs sparsity strategies (e.g., rule-based rewards or information density scoring) to filter redundant visual or semantic features, while the slow path focuses on refined reasoning.
-This hierarchical mechanism not only reduces computational overhead by up to 30-50%, but also continuously optimizes token selection strategies via RL’s online feedback.
-The sparsity reinforcement paradigm proposed in ZipR1 chen2025zipr1reinforcingtokensparsity highlights RL’s unique advantage in cross-modal generalization.
-By designing sparsity rewards (e.g., token quantity limits or step compression incentives), models can autonomously generate concise intermediate reasoning chains and even distill such capabilities into lightweight models.
-For example, in visual question answering tasks, models can dynamically extract key image region features while enhancing answer reliability through multi-candidate contrast mechanisms (e.g., parallel samplings), achieving the goal of “low computational cost, high task accuracy.”
-Recent findings further reveal that optimizing only a small subset of high-entropy "forking tokens" can match or even surpass full-gradient updates in reasoning tasks wang2025beyond , suggesting that token-level entropy patterns can guide efficient policy updates.
-This entropy-aware strategy reinforces the broader view that sparse, selective token refinement is sufficient for preserving reasoning performance in large models.
-Looking ahead, the integration of such approaches could enhance cross-modal alignment and inference efficiency in real-time and resource-constrained scenarios, supporting a new generation of lightweight yet capable multimodal language agents.
+5.7.
 5.2 From Prompt Tuning to Chain of Thought Reasoning
-Current token reduction efforts for prompts have primarily aimed at compressing prompts for efficiency, often with impressive results jiang2023llmlingua ; mu2023learning .
+Current token reduction efforts for prompts have primarily aimed at compressing prompts for efficiency, often with impressive results Jiang et al.
+(2023); Mu et al.
+(2023).
 Looking forward, token reduction should evolve into enhancing reasoning and maximizing utility per token in context.
 Instead of focusing solely on making prompts shorter, future research should explore how each remaining token can carry more information or trigger more complex inference during in-context learning and chain of thought reasoning.
-One direction is to alter the generation paradigm itself, for example, training language models to predict multiple tokens per step gloeckle2024better .
-Another idea is to enable deeper internal reasoning without increasing prompt length adams2023sparse .
+One direction is to alter the generation paradigm itself, for example, training language models to predict multiple tokens per step Gloeckle et al.
+(2024).
+Another idea is to enable deeper internal reasoning without increasing prompt length Adams et al.
+(2023).
 As mentioned in Sec.
-4, long CoT chains can become verbose: excessive reasoning steps may introduce errors or obscure logical clarity, particularly in LLM agents where internal reasoning alternates with external tool use gao2025txagent ; wang2024toolgen .
+4, long CoT chains can become verbose: excessive reasoning steps may introduce errors or obscure logical clarity, particularly in LLM agents where internal reasoning alternates with external tool use Gao et al.
+(2025a); Wang et al.
+(2024b).
 Token reduction may serve a critical role in this context by compressing intermediate reasoning into a compact representation.
-For example, approaches based on next-token prediction mu2023learning ; zhang2025lightthinker can distill intermediate thinking chains into a set of dense, information-rich tokens.
+For example, approaches based on next-token prediction Mu et al.
+(2023); Zhang et al.
+(2025b) can distill intermediate thinking chains into a set of dense, information-rich tokens.
 These compressed representations can then replace the full intermediate context and serve as inputs for subsequent reasoning steps.
@@
 This involves training models with objectives that reward higher-level inference per token, developing architectures that recycle tokens for multi-step thinking, or dynamically selecting the most salient tokens to keep at each step of reasoning.
+5.3 Efficient Reasoning with Reinforcement Learning
+Reinforcement-learning (RL)-driven token reduction has shown strong promise for improving reasoning efficiency in both LLMs and MLLMs.
+The key challenge is to balance compute and reasoning quality via dynamic reward design, sparsity-inducing constraints, and adaptive control of effective token length Aggarwal and Welleck (2025); Luo et al.
+(2025); Wu et al.
+(2025c).
+In language reasoning, length awareness is incorporated either by diversifying reasoning formats during pre-training Su et al.
+(2025a) or by adding explicit length penalties in the RL stage Arora and Zanette (2025); Hou et al.
+(2025).
+While effective, most approaches implicitly assume static task complexity or rely on hand-specified length constraints, which can be suboptimal under heterogeneous workloads.
+A natural next step is adaptive reasoning: using RL to learn per-instance budget allocation from intrinsic task difficulty NVIDIA (2025); Qwen (2025).
+In parallel, performing reasoning in a compressed latent space can yield substantial computational savings Shen et al.
+(2025b), but current methods often degrade due to poorly structured latent representations.
+A promising direction is to inject explicit logical structures into the latent space to enable more controllable and compositional reasoning Helff et al.
+(2025).
+Beyond language modality, under the “Fast-Slow Thinking” framework Xiao et al.
+(2025a), RL can supervise hierarchical selection of high-value tokens: a fast branch applies sparsity signals (e.g., rule-based rewards or information-density scoring) to prune redundant visual/semantic features, while a slow branch allocates computation to refined reasoning.
+Additionally, RL enables a Think-with-Image paradigm Su et al.
+(2025d) by letting models adapt visual granularity: VisionThink Yang et al.
+(2025b) adopts a progressive resolution strategy, using RL to selectively request high-resolution inputs only for fine-grained tasks like OCR.
+PixelThink Wang et al.
+(2025c) addresses visual overthinking in segmentation by adjusting reasoning length based on task difficulty.
+These methods demonstrate how RL can dynamically calibrate both input resolution and reasoning depth across various tasks.
+Looking ahead, the integration of such approaches could enhance cross-modal alignment and inference efficiency in real-time and resource-constrained scenarios, supporting a new generation of lightweight yet capable multimodal language agents.
 5.4 Complementary to Other Methods
 Token reduction can complement other efficiency techniques, such as quantization.
-By selectively reducing the number of tokens processed during inference, models can improve both performance and efficiency, particularly when paired with quantization strategies li2025mergevq .
+By selectively reducing the number of tokens processed during inference, models can improve both performance and efficiency, particularly when paired with quantization strategies Li et al.
+(2025b).
 Traditional key-value cache quantization methods often suffer from accuracy loss due to their inability to handle outlier tokens that carry distinct or rare features.
-To mitigate this issue, Outlier Token Tracking su2025accuratekvcachequantization identifies outlier tokens during decoding and excludes them from quantization, preserving full-precision necessary representations and improving key-value cache quantization accuracy.
-Similarly, Agile-Quant shen2024agile incorporates token pruning as a preprocessing step to reduce the impact of activation outliers.
+To mitigate this issue, Outlier Token Tracking Su et al.
+(2025c) identifies outlier tokens during decoding and excludes them from quantization, preserving full-precision necessary representations and improving key-value cache quantization accuracy.
+Similarly, Agile-Quant Shen et al.
+(2024) incorporates token pruning as a preprocessing step to reduce the impact of activation outliers.
 It prunes tokens based on their attention to the start-of-sequence token, discarding those with low attentiveness, which often appear in adjacent channels and contribute to quantization noise.
@@
 5.5 Towards Dense Prediction Tasks for Vision
-Existing works primarily concentrate on compressing the backbone of models to ensure their generalization ability, and few works explore recovering all tokens for dense prediction tasks bolya2023token ; liang2022expediting .
-It is necessary to develop custom token reduction methods for various downstream dense prediction applications like autonomous driving and robotic control with specific settings and requirements.
+Existing works primarily concentrate on compressing the backbone of models to ensure their generalization ability, and few works explore recovering all tokens for dense prediction tasks Bolya and Hoffman (2023); Liang et al.
+(2022a).
+It is necessary to develop custom token reduction methods for various downstream dense prediction applications like autonomous driving and robotic control with specific settings and requirements Cao et al.
+(2025); Pei et al.
+(2025).
 Lacking these specialized designs would lead to a mismatch and performance drop when deployed in real-world settings.
-For example, autonomous driving zhang2024sparsead would require displacement and velocity based on occupancy prediction, and robotic control wang2024sparse would demand rotation angle according to the grid map.
+For example, autonomous driving Zhang et al.
+(2024a) would require displacement and velocity based on occupancy prediction, and robotic control Wang et al.
+(2024c) would demand rotation angle according to the grid map.
 Therefore, how to develop fast and specialized token reduction strategies tailored for downstream dense prediction tasks is crucial for deployment in practical scenarios.
@@
 Exploiting long videos holds great potential, as processing hours of footage is significantly more labor-intensive and time-consuming than working with short clips.
-Due to the inherent complexity and resource demands, most current research on long video learning focuses on discriminative tasks such as video understanding lee2024video ; ren2023testa .
-In contrast, broader applications including long video editing zhang2025adaflow , long video-text retrieval, and narrative-level generation remain largely underexplored.
+Due to the inherent complexity and resource demands, most current research on long video learning focuses on discriminative tasks such as video understanding Lee et al.
+(2024); Ren et al.
+(2023).
+In contrast, broader applications including long video editing Zhang et al.
+(2025f), long video-text retrieval, and narrative-level generation remain largely underexplored.
 Progress in these areas could have a significant impact on scene editing in video clips, character rendering in movies, and retrieving useful information from numerous videos.
-Moreover, token reduction offers a path toward interpretability and efficiency in long video processing jiang2025token .
+Moreover, token reduction offers a path toward interpretability and efficiency in long video processing Jiang et al.
+(2025a).
 This mimics the human visual system, which does not attend to every frame in detail but instead focuses on salient spatiotemporal changes, such as actions or object movement, while filtering out static, redundant content like backgrounds and stationary objects.
 Future models should similarly prioritize informative frames and temporal segments, allowing them to reason over extended video sequences with greater efficiency and interpretability.
+5.7 Algorithm-Hardware Co-Design
+While algorithmic advancements in token reduction have achieved impressive computational savings, the next crucial step is to integrate these techniques with hardware-aware design principles.
+We posit that algorithm-hardware co-design is essential for holistic optimization across the compute stack, considering the interplay between algorithmic choices, hardware architectures (specialized data paths, memory hierarchies, communication fabrics, control logic, etc.), and compiler/runtime support (efficient sparse mapping, dynamic scheduling, irregular-data management, etc.) Dong et al.
+(2023); Parikh et al.
+(2024).
+Currently, co-design efforts targeted at token reduction lag significantly behind pure algorithmic research.
+This gap is problematic because hardware design needs to balance PPA (power, performance, and area), platform specifics, data movement costs, control overhead, and scalability/reusability Pan et al.
+(2024).
+Algorithms developed in isolation often generate sparse or irregular compute patterns that general-purpose hardware cannot exploit effectively.
+Therefore, future research should aim to: 1) Design parameterizable, reconfigurable accelerator modules-such as on-the-fly importance-scoring units and sparse-data pipelines-that natively support token-reduced Transformers.
+2) Explore Processing-in-Memory (PIM) architectures to alleviate severe memory bottlenecks caused by dynamic token pruning.
+By executing scoring operations or partial attention mechanisms within or near memory arrays, PIM can drastically reduce data movement costs and improve end-to-end efficiency.
+5.8 Towards Efficient Agentic Systems
+Dynamic Memory & Context Engineering.
+AI agents face a severe context bottleneck where accumulating interaction history not only incurs quadratic computational costs but also degrades reasoning through "lost-in-the-middle" phenomena.
+Transitioning to active context management is essential Anthropic (2025); systems must distinguish between immutable instructions and transient episodic data.
+ACON Kang et al.
+(2025) demonstrate that semantic memory compression can reduce peak token usage while maintaining long-horizon performance.
+Complementary strategies include hierarchical memory abstraction, which offloads older history to retrieval-based storage, and hybrid masking techniques JetBrains Research (2025).
+These approaches treat tokens as a finite resource to be optimized on-the-fly, dynamically adjusting the sparsity of the context window based on the complexity of the current reasoning step.
+Observation & Tool Pruning.
+Agents must often process verbose tool outputs (e.g., massive JSON/HTML).
+Feeding raw data is inefficient; future research should focus on token-aware interaction, where lightweight scorers filter observations to retain only task-relevant features.
+This preserves critical context bandwidth without sacrificing trajectory accuracy, as evidenced in recent programming agents Xiao et al.
+(2025b).
+Furthermore, adaptive truncation strategies that prioritize error traces over standard success logs can significantly improve debugging capabilities while minimizing token consumption.
+Communication-Efficient Multi-Agent Systems.
+At the system level, the aggregate token cost scales with the number of interacting nodes.
+Enforcing strict token budgets on inter-agent exchange compels information-dense communication.
+This sparse protocol enhances scalability by reducing redundant chatter and mitigating hallucination loops in collaborative workflows Zhang et al.
+(2024b); Zou et al.
+(2025b).
 5.9 Towards AI for Broader ML and Scientific Domains
@@
 Building Biomedical Tokenizers.
-Recent works exemplify the transformative potential of advanced tokenization methods in the biomedical domain, including protein yuan2025protein ; gao2025foldtoken ; suyunu2025evobpe , genomic eapen2025genomic , and chemical structure yan2024invariant tokenizers.
+Recent works exemplify the transformative potential of advanced tokenization methods in the biomedical domain, including protein Gao et al.
+(2025b); Suyunu et al.
+(2025); Yuan et al.
+(2025), genomic Eapen (2025), and chemical structure Yan et al.
+(2024) tokenizers.
 Collectively, these methods illustrate how informed reduction and condensation of input tokens can lead to more effective and interpretable scientific models.
 For example, traditional tokenizers in EHR foundation models typically treat medical codes as isolated textual units, neglecting their inherent structured and relational context, such as hierarchical relationships, disease co-occurrences, and drug-treatment associations found within biomedical ontologies.
-To solve this issue, MedTok su2025multimodal integrates textual descriptions and graph-based relational data into a unified tokenization framework.
+To solve this issue, MedTok Su et al.
+(2025b) integrates textual descriptions and graph-based relational data into a unified tokenization framework.
 It first uses a language model encoder to extract embeddings from medical code descriptions and employs a graph encoder to capture relational structures from biomedical ontologies.
@@
 This process drastically reduces effective vocabulary size, addressing the scalability challenge of 600,000+ medical codes by collapsing redundant representations while preserving critical clinical context.
-Inspired by adaptive tokenization methods for vision duggal2024adaptive ; kang2023tictok ; yan2024elastictok , future EHR tokenization would be adaptive, enabling the dynamic representation of patients’ medical histories, where the length of the token series for each patient’s history would be directly correlated with the length and complexity.
+Inspired by adaptive tokenization methods for vision Duggal et al.
+(2024); Kang and Lee (2023); Yan et al.
+(2025), future EHR tokenization would be adaptive, enabling the dynamic representation of patients’ medical histories, where the length of the token series for each patient’s history would be directly correlated with the length and complexity.
 Such adaptive tokenization can significantly improve training and inference efficiency across diverse healthcare systems.
@@
 Temporal dynamics form an essential component of clinical reasoning, particularly through longitudinal patient data like lab tests and vital signs.
-However, current large language models struggle to effectively incorporate time-series inputs due to challenges in temporal tokenization spathis2024first ; anjum2024lipcot ; fang2025tsla ; spathis2024first ; masserano2024enhancing .
-Future tokenization methods should not only dynamically adjust the number of tokens according to temporal complexity but also selectively focus on time segments most relevant to the clinical context, prompt, or task at hand talukder2024totem .
+However, current large language models struggle to effectively incorporate time-series inputs due to challenges in temporal tokenization Spathis and Kawsar (2024); Anjum (2024); Fang et al.
+(2025b); Spathis and Kawsar (2024); Masserano et al.
+(2024).
+Future tokenization methods should not only dynamically adjust the number of tokens according to temporal complexity but also selectively focus on time segments most relevant to the clinical context, prompt, or task at hand Talukder et al.
+(2024).
 This could enhance training effectiveness and inference accuracy, helping create the next generation of EHR foundation models, which are flexible not only over different tasks or prompts, but also over different data sources, patients, and populations.
@@
 They could facilitate timely disease predictions, accurately forecast chronic disease trajectories, and anticipate patient responses to treatments.
-5.8 Algorithm-Hardware Co-Design
-While algorithmic advancements in token reduction have achieved impressive computational savings, the next crucial step is to integrate these techniques with hardware-aware design principles.
-We posit that algorithm-hardware co-design is essential for holistic optimization across the compute stack, considering the interplay between algorithmic choices, hardware architectures (specialized data paths, memory hierarchies, communication fabrics, control logic, etc.), and compiler/runtime support (efficient sparse mapping, dynamic scheduling, irregular-data management, etc.) dong2023heatvit ; parikh2024accelerating .
-Currently, co-design efforts targeted at token reduction lag significantly behind pure algorithmic research.
-This gap is problematic because hardware design needs to balance PPA (power, performance, and area), platform specifics, data movement costs, control overhead, and scalability/reusability 10.1109/ASP-DAC58780.2024.10473968 .
-Algorithms developed in isolation often generate sparse or irregular compute patterns that general-purpose hardware cannot exploit effectively.
-Therefore, future research should aim to: 1) Design parameterizable, reconfigurable accelerator modules-such as on-the-fly importance-scoring units and sparse-data pipelines-that natively support token-reduced Transformers.
-2) Explore Processing-in-Memory (PIM) architectures to alleviate severe memory bottlenecks caused by dynamic token pruning.
-By executing scoring operations or partial attention mechanisms within or near memory arrays, PIM can drastically reduce data movement costs and improve end-to-end efficiency.
 6 Conclusion
@@
 We have shown how principled token reduction can address key challenges such as enhancing semantic fidelity in vision-language alignment, curbing verbose reasoning trajectories, preserving long-range coherence, and stabilizing learning dynamics.
-Looking ahead, we outlined a roadmap of promising directions.
-We anticipate that embracing token reduction as a holistic, task-aware mechanism will yield the next generation of generative architectures that are not only more efficient but also more robust, interpretable, and aligned with real-world demands.
+Looking forward, the roadmap we outlined points to a broad landscape of opportunities, ranging from algorithmic innovations and hardware-algorithm co-design to specialized applications in scientific domains.
+We anticipate that future work will increasingly focus on constructive compression and reinforcement learning-guided selection, enabling models to autonomously optimize their information bandwidth.
+Ultimately, by treating token reduction as a holistic and task-aware mechanism, the community can develop next-generation systems that effectively balance scalability with effectiveness, interpretability, and performance.
+7 Limitations
+While token reduction offers significant benefits, our review identifies critical limitations and trade-offs that must be considered to avoid indiscriminate application.
+Information Loss in Dense Prediction
+Token reduction methods, particularly pruning, inherently discard information.
+While this is acceptable for semantic classification or generation, it poses severe risks for dense prediction tasks (e.g., segmentation, object detection) or medical analysis where fine-grained spatial details are crucial.
+Merging strategies like ToMe Bolya and Hoffman (2023) mitigate this better than pruning, but artifacts often remain at high compression ratios.
+In scenarios requiring pixel-perfect reconstruction, the trade-off between reduction and precision often favors preserving the full token set.
+Furthermore, the lack of specialized designs for token recovery often leads to performance mismatches in real-world applications like autonomous driving and robotic control.
+Future research must therefore explore custom reconstruction mechanisms to ensure these methods can meet the rigorous demands of dense tasks.
+Overhead vs. Gain
+Dynamic token reduction introduces computational overhead (e.g., scoring networks, predictors).
+For short sequences or small batch sizes, the cost of computing token importance may outweigh the savings from processing fewer tokens.
+Furthermore, unstructured pruning can lead to irregular memory access patterns that are inefficient on standard hardware (GPUs/TPUs), potentially negating theoretical FLOPs reductions.
+Alternatives: Reduction vs. Retrieval
+Critics may argue that techniques like Retrieval-Augmented Generation (RAG) or simply scaling context windows render token reduction unnecessary.
+However, we argue these are complementary.
+While RAG selects documents, token reduction operates at the sub-document level, filtering noise within relevant chunks.
+Similarly, while larger context windows allow for more input, they exacerbate the "lost-in-the-middle" phenomenon; token reduction acts as an attention-sharpening mechanism that helps models utilize these long contexts more effectively.
+Appendix A Theoretical Formulation
+In this section, we provide a unified mathematical framework for token reduction methods.
+We formalize the token reduction process into two distinct phases: Compression Criteria (how to evaluate tokens) and Compression Strategies (how to reduce tokens).
+A.1 Problem Definition
+Given an input sequence $X=[x_{1},x_{2},\dots,x_{N}]\in\mathbb{R}^{N\times D}$, where $N$ represents the sequence length and $D$ denotes the feature dimension, the objective of token reduction is to generate a compressed sequence $X^{\prime}=[x^{\prime}_{1},x^{\prime}_{2},\dots,x^{\prime}_{M}]\in\mathbb{R}^{M\times D}$, such that $M\ll N$.
+The reduction process can be formally defined as a composite function of a scoring criterion $\mathcal{E}$ and a compression strategy $\mathcal{P}$:
+$X^{\prime}=\mathcal{P}(X,\mathcal{E}(X))$ (1)
+where $\mathcal{E}(X)$ outputs importance scores, clustering assignments, or gradient sensitivities, and $\mathcal{P}$ executes the dimensionality reduction.
+Figure 2 shows the token reduction pipeline.
+A.2 Compression Criteria
+The scoring function $\mathcal{E}:X\to\mathcal{S}$ determines the semantic value or redundancy of each token.
+We broaden the categorization to include gradient and entropy-based metrics alongside standard parametric/non-parametric approaches.
+Attention-based Scoring.
+Utilizing the inherent sparsity of the self-attention mechanism, the importance of a token $x_{i}$ is quantified by the attention it receives.
+This can be global (averaged across all heads/tokens) or targeted (attention from the special [CLS] token or specific query tokens).
+The score $s_{i}$ is typically calculated as:
+$s_{i}=\sum_{j\in\mathcal{Q}}\text{Attn}(x_{j},x_{i})$ (2)
+where $\mathcal{Q}$ is the set of query tokens (e.g., $\mathcal{Q}=\{x_{\text{CLS}}\}$ for classification tasks or $\mathcal{Q}=\{x_{1\dots N}\}$ for global density).
+Similarity-based Scoring.
+This approach assumes that tokens close in the feature space contain redundant information.
+The criterion calculates pairwise distances to identify clusters or redundant pairs.
+For tokens $(x_{i},x_{j})$, the metric is typically cosine similarity:
+$\text{Sim}(x_{i},x_{j})=\frac{x_{i}\cdot x_{j}}{\|x_{i}\|_{2}\|x_{j}\|_{2}}$ (3)
+High similarity scores ($\text{Sim}>\tau$) trigger merging operations. Advanced methods extend this to density-based clustering (e.g., K-Means or DPC-KNN) to identify representative centroids.
+Gradient and Entropy-based Scoring.
+Beyond static feature analysis, recent methods employ dynamic metrics.
+Gradient-based criteria measure a token’s contribution to the loss function, retaining tokens with high gradient norms ($\|\nabla_{x_{i}}\mathcal{L}\|$).
+Entropy-based criteria evaluate the uncertainty of the model’s prediction; tokens with low information density (low entropy) are candidates for pruning in early-exit or fast-forwarding frameworks.
+Parametric Scoring. Parametric methods introduce a lightweight auxiliary module (e.g., a predictor network $\mathcal{M}_{\phi}$) to explicitly predict token utility:
+$S=\mathcal{M}_{\phi}(X)$ (4)
+where $S\in[0,1]^{N}$ represents the keep-probability or saliency score.
+These predictors are trained via Gumbel-Softmax or reinforcement learning (RL) to maximize downstream accuracy while minimizing token count.
+A.3 Compression Strategies
+[…]
+Figure 2: The token reduction pipeline. We formulate reduction as a composite of Criteria $\mathcal{E}$ (scoring) and Strategy $\mathcal{P}$ (pruning/merging).
+Once the relationships or scores are established, the compression strategy $\mathcal{P}$ transforms the sequence. We classify these into four primary mechanisms.
+Token Pruning (Hard & Soft). Pruning is a selection process that discards tokens based on the criteria $\mathcal{E}$.
+$X^{\prime}=\{x_{i}\mid s_{i}\in\text{TopK}(S)\lor s_{i}>\tau\}$ (5)
+While Hard Pruning permanently removes tokens, Soft Pruning (or "packaging") aggregates the discarded tokens into a single summary token to preserve residual information, preventing total information loss.
+Token Merging & Clustering.
+Merging aggregates information from a set of tokens $\mathcal{C}=\{x_{1},\dots,x_{k}\}$ identified as similar.
+This ranges from bipartite matching (merging pairs) to density-based clustering (merging large groups).
+The merged token $x^{\prime}_{\text{cluster}}$ is a weighted average:
+$x^{\prime}_{\text{cluster}}=\frac{\sum_{x_{j}\in\mathcal{C}}w_{j}x_{j}}{\sum_{x_{j}\in\mathcal{C}}w_{j}}$ (6)
+where $w_{j}$ tracks the token’s size (number of constituent patches), ensuring proportional representation of fused features.
+Transformation-based Compression. These methods reduce sequence length through structural operations, exploiting spatial (image) or temporal (video) priors. Common techniques include:
+  • Pooling/Unshuffle: Non-parametric downsampling: $\text{Pool}:\mathbb{R}^{H\times W}\to\mathbb{R}^{\frac{H}{r}\times\frac{W}{r}}$.
+  • Convolution: Strided convolution to abstract local neighborhoods: $X^{\prime}=\text{Conv}_{k\times k}(X,s)$.
+Token Distillation (Query-based).
+Distillation employs a set of learnable latent queries $Q\in\mathbb{R}^{M\times D}$ to extract information from the input $X$ via cross-attention mechanisms (e.g., Perceiver Resampler or Q-Former).
+This decouples output length $M$ from input length $N$:
+$X^{\prime}=\text{Softmax}\left(\frac{Q(XW_{K})^{T}}{\sqrt{D}}\right)(XW_{V})$ (7)
+This strategy is particularly effective for cross-modal alignment, compressing dense visual features into sparse text-aligned tokens.
+[…]
+Figure 3: Visualization of token reduction.
+(a) Image: Visual tokens are pruned based on saliency, retaining only the most salient patches.
+(b) Text: Low-information stop words (gray) are removed to form a compressed semantic core.
+We summarize the general workflow of training-free token reduction in Algorithm 1. We also show a visualization of token reduction in Figure 3.
+Algorithm 1 General Training-Free Token Reduction Workflow
+1:Input tokens $X\in\mathbb{R}^{N\times D}$, Target ratio $r$
+2:Compressed tokens $X^{\prime}\in\mathbb{R}^{M\times D}$
+[…]
+4:Phase 1: Criteria Calculation ($\mathcal{E}$)
+5:if Attention-based then
+6:  $S\leftarrow\text{Agg}(\text{AttentionMap}(X))$ $\triangleright$ Agg: Sum/Avg over heads
+7:else if Similarity-based then
+8:  $A\leftarrow XX^{T}/(\|X\|\|X\|)$ $\triangleright$ Cosine Similarity Matrix
+9:  Partition $X$ into sets $\{\mathcal{C}_{1},\dots,\mathcal{C}_{M}\}$ via clustering on AA
+10:end if
+11:Phase 2: Strategy Execution ($\mathcal{P}$)
+12:if Pruning then
+13:  $\text{Indices}\leftarrow\text{TopK}(S,M)$
+14:  $X^{\prime}\leftarrow\text{Gather}(X,\text{Indices})$
+15:else if Merging then
+16:  for $m=1$ to $M$ do
+17:   $x^{\prime}_{m}\leftarrow\text{WeightedSum}(\mathcal{C}_{m})$
+18:  end for
+19:end if
+20:return $X^{\prime}$
+A.4 Controllable Reasoning via Reinforcement Learning
+Controllable Reasoning refers to the ability of language models to dynamically adjust the depth and length of their reasoning processes according to user-specified constraints (e.g., exact or maximum token lengths), thereby enabling a tunable trade-off between reasoning efficiency and accuracy.
+State-of-the-art approaches Luo et al.
+(2025); Aggarwal and Welleck (2025) typically employ reinforcement learning frameworks Shao et al.
+(2024), where a multi-objective reward function is designed to optimize correctness and length compliance jointly.
+Specifically, the reward function incorporates two key objectives:
+  1. 1.
+    Correctness reward: awarded if the model’s output matches the ground-truth answer;
+  2. 2.
+    Length penalty: imposed if the generated sequence length deviates from the target length.
+Exact Length Control. It requires the model to generate reasoning sequences whose length exactly matches a user-specified target length:
+$R_{\text{exact}}(y,y_{\text{t}},n_{\text{t}})=\mathbb{I}(y=y_{\text{t}})-\alpha\cdot|n_{\text{t}}-n_{y}|,$ (8)
+where $y$ is the generated sequence, $y_{\text{t}}$ is the ground truth answer, $n_{\text{t}}$ is the target token length, $n_{y}$ is the actual token length of the generated sequence, $\mathbb{I}(\cdot)$ is the indicator function (1 if correct, 0 otherwise), and $\alpha$ is a penalty weight balancing correctness and length.
+Maximum Length Control. It controls the model to generate reasoning sequences no longer than a specified upper limit, encouraging efficient reasoning within a token budget:
+$R_{\text{max}}(y,y_{\text{t}},n_{\text{t}})=\mathbb{I}(y=y_{\text{t}})\cdot\text{clip}\left(\alpha\cdot(n_{\text{t}}-n_{y})+\delta,0,1\right),$ (9)
+where $\text{clip}(\cdot,0,1)$ clamps reward to the range $[0,1]$ and $\delta$ is an offset term to avoid zero reward (typically set to 0.5).
+Length Efficiency Optimization. It encourages the model to shorten reasoning length while maintaining correctness, particularly useful for reducing redundancy in long-reasoning models:
+$R_{\text{eff}}(y,y_{\text{t}},x)=\frac{\bar{L}_{\text{ref}}(x)}{L(y)}-1+\lambda\left(A(y,y_{\text{t}})-\bar{A}_{\text{ref}}(x)\right),$ (10)
+where $\bar{L}_{\text{ref}}(x)$ is the average length of reference model outputs for problem $x$, $A(y,y_{\text{t}})$ is the accuracy function (1 if correct, 0 otherwise), $\bar{A}_{\text{ref}}(x)$ is the average accuracy of the reference model on problem $x$, and $\lambda$ is an accuracy penalty weight to prevent performance degradation from over-compression.
+By adjusting hyperparameters (e.g., $\alpha$ and $\lambda$) in the reward function, one can flexibly control the model’s tendency to prioritize correctness versus length, which not only achieves precise length control but also maintains or even improves model performance while significantly reducing reasoning overhead.